### v2

# Concept Unlearning by Modeling Key Steps of Diffusion Process

Abstract
Text-to-image diffusion models (T2I DMs), represented by Stable Diffusion, which generate highly realistic images based on textual input, have been widely used. However, their misuse poses serious security risks. While existing concept unlearning methods aim to mitigate these risks, they struggle to balance unlearning effectiveness with generative retainability.To overcome this limitation, we innovatively propose the Key Step Concept Unlearning (KSCU) method, which ingeniously capitalizes on the unique stepwise sampling characteristic inherent in diffusion models during the image generation process. Unlike conventional approaches that treat all denoising steps equally, KSCU strategically focuses on pivotal steps with the most influence over the final outcome by dividing key steps for different concept unlearning tasks and fine-tuning the model only at those steps. This targeted approach reduces the number of parameter updates needed for effective unlearning, while maximizing the retention of the model’s generative capabilities.Through extensive benchmark experiments, we demonstrate that KSCU effectively prevents T2I DMs from generating undesirable images while better retaining the model’s generative capabilities. Our code will be released.
Preprint. Under review.
Figure 1: Compared to previous methods (b), KSCU (c) focuses exclusively on the denoising steps that have the most significant impact on concept generation and achieves effective forgetting of the target concept while better preserving the generation of unrelated concepts.
1 Introduction
In recent years, the advancement of text-to-image (T2I) diffusion models [22, 20, 9, 16, 26, 43] has led to their widespread adoption across various domains [38, 31, 42, 1, 27], including short videos, comics, and illustrations. As a result, AI-generated images have become an integral part of daily life for many individuals. However, large-scale T2I models are trained on extensive datasets [45, 4, 37, 38] that often contain sensitive content, such as unauthorized artistic works [46, 47] or Not Safe For Work (NSFW) material. Consequently, these models can inadvertently regenerate infringing content when prompted with relevant text inputs [3]. This raises significant concerns regarding copyright infringement and social security risks, highlighting the urgent need for effective mitigation strategies.
The generation of NSFW content poses a major security risk for mainstream text-to-image diffusion models. Researchers have explored various approaches to mitigate this issue, including dataset-based screening, content filtering [39], and fine-tuning techniques [7, 30, 41]. concept unlearning, as a fine-tuning-based approach [23, 11, 56, 13, 14, 53, 18], selectively adjusts partial model parameters to remove the model’s ability to generate specific concepts. Through concept unlearning, T2I diffusion models effectively ‘forget’ designated visual concepts by replacing them with semantically neutral content. This targeted parameter adjustment (referred to as ”concept erase” [12, 21, 14]) provides distinct advantages over conventional methods—offering greater robustness than post-hoc filtering [39] while avoiding the intensive labor of dataset curation [40], and has become one of the most promising approaches to curb the generation of NSFW content in T2I propagation models.
Although significant progress has been made in concept unlearning for T2I diffusion models, several challenges remain unresolved. a. Blind unlearning of the entire diffusion process. Existing concept unlearning approaches for T2I diffusion models fail to notice for the stepwise denoising characteristics of diffusion models. Since diffusion models generate images through a progressive denoising process—where earlier steps primarily capture low-frequency structural information and later steps refine high-frequency details [17, 50, 55, 35, 2]—unlearning does not necessarily require fine-tuning across all denoising steps. b. Unbalanced unlearning effectiveness and generative retainability. Most existing methods require extensive modifications to the original model parameters to ensure unlearning effectiveness. However, excessive parameter updates may lead to over-unlearning, negatively affecting the model’s ability to generate unrelated concepts. Conversely, strictly limiting the scale of parameter updates to preserve generative retainability—i.e., the model’s ability to retain generation quality for unrelated concepts—makes it difficult to fully unlearn the target concept. c. Robustness at high cost. Adversarial attack [51, 6, 58] research on unlearned models has demonstrated that carefully crafted adversarial text prompts can successfully recover unlearned concepts. While adversarial training [21, 59] can enhance robustness against such attacks, it is computationally expensive and brings low generative retainability.
To address existing challenges, we propose Key Step Concept Unlearning (KSCU), a novel framework for concept unlearning in T2I diffusion models. Unlike prior methods that operate across all denoising steps, KSCU selectively fine-tunes a subset of key steps based on a predefined Key Step Table. For different unlearning tasks, multiple step intervals are empirically defined, and a selection algorithm prioritizes steps with the greatest influence on the final output, enabling efficient and targeted unlearning. To improve robustness, KSCU applies Prompt Augmentation in the final training stage, including random shuffling, word removal, character insertion, and embedding noise. These augmented prompts help the model better capture the semantic distribution of the unlearned concept in the embedding space. Additionally, we introduce a Key Step Unlearning Loss tailored to classifier-free guidance DMs. By combining Key Step Table, Prompt Augmentation, and Key Step Unlearning Loss, KSCU achieves more effective concept unlearning while better ensuring generative retainability.
Our contributions can be summarized as follows:
  • We propose KSCU, a novel concept unlearning framework for T2I DMs. Unlike prior work, KSCU only performs fine-tuning for the key step, achieving efficient unlearning by applying smaller adjustments to the model, thereby maintaining the model’s generation ability, taking both unlearning effectiveness and generative retainability into account.
  • Our proposed KSCU integrates three innovations: Key Step Table, a step selection strategy focusing on critical denoising steps to achieve effective concept unlearning with minimal fine-tuning overhead; Prompt Augmentation to enhance robustness by introducing diverse prompt variations; and Key Step Unlearning Loss, a concept unlearning loss tailored for classifier-free guided diffusion models, enabling effective concept removal while preserving generation quality.
  • Extensive experimental results on multiple datasets demonstrate the superiority of KSCU. For instance, in the NSFW unlearning experiment, KSCU achieves a remarkable unlearning accuracy of 96.5% with FID of just 14.1, significantly outperforming the previous SOTAs.
2 Related Work
2.1 Text-to-Image Diffusion Models
Diffusion models [17] were initially proposed as a generative framework based on the reverse denoising process to reconstruct target input images. With the introduction of advanced techniques such as score-based models [50], noise-conditioned score network [49], and denoising-based guidance [9], these models have become increasingly popular for text-to-image (T2I) generation. The emergence of latent diffusion models (LDMs) [40] and classifier-free guidance [16] has further improved the controllability and efficiency of diffusion-based T2I systems. Their ability to generate high-quality, semantically coherent images has established diffusion models as the dominant approach in T2I generation.
Building on these advancements, recent research has focused on optimizing T2I diffusion models in terms of efficiency, scalability, and applicability. Significant breakthroughs have been made in acceleration algorithms [48, 25, 28, 29], model architectures [40], and large-scale datasets [45, 4, 37, 38], all of which have contributed to the widespread adoption of these models in practical applications [38, 31, 42, 1]. However, despite their advantages, these models are vulnerable to misuse, leading to concerns over NSFW content [39], unauthorized depictions of individuals, and copyright-infringing artwork [1]. As T2I diffusion models continue to advance in realism and fidelity, addressing their ethical and legal challenges has become an urgent priority for the research community.
Figure 2: The KSCU framework consists of three key modules: Key Step Unlearning Loss, Key Step Table, and Prompt Augmentation. The Key Step Table predefines step selection strategies for four concept unlearning tasks (Class, Style, NSFW, and Instance). During training, a denoising step $t$ is selected from the Key Step Table. The unlearn model performs denoising sampling to obtain $x_{t}$, while the Prompt Augmentation module converts the target concept $c$ into an augmented prompt $c^{*}$. Using $x_{t}$, $t$, $c$, and $c^{*}$, the Key Step Unlearning Loss is computed via Equation (7), and designated parameters of the unlearn model are updated through backpropagation.
2.2 Diffusion Model Concept Unlearning
Large-scale datasets [45, 4, 37, 38] often contain undesirable data, including content with inherent biases, ethically problematic information, or copyrighted materials that may later become restricted. Models trained on such datasets acquire the ability to generate this sensitive content, posing significant security and ethical risks in text-to-image (T2I) generation. To mitigate these concerns, concept unlearning in diffusion models has been proposed, aiming to selectively erase specific generative capabilities while preserving the model’s ability to generate unrelated content—without requiring full retraining.
Several existing methods have been developed for concept unlearning in T2I diffusion models. ESD [12] and CA [23] employ a frozen guiding model to provide negative guidance for fine-tuning. UCE [13] and RECE [14] leverages a closed-form solution to modify cross-attention weights, enabling simultaneous unlearning of multiple concepts. Salun [11] and Scissorhands [52] optimize parameter updates based on their influence on concept retention. Meanwhile, RACE [21] and AdvUnlearn [59]generate adversarial samples to enhance unlearning and improve model robustness.
Although concept unlearning methods provide a certain level of security for T2I diffusion models, their effectiveness remains questionable when faced with diverse real-world scenarios and adversarial examples. Recent studies on adversarial attacks [51, 6, 58] against T2I DMs indicate that existing concept unlearning methods remain vulnerable to such attacks. To enhance robustness, current approaches rely on increasing the number of fine-tuning iterations or incorporating adversarial training [59, 21]. However, these methods significantly increase computational cost while further degrading the model’s generative quality.
Unlike previous approaches [23, 12, 53, 33], KSCU explicitly considers the relationship between the stepwise denoising process of diffusion models and the effectiveness of concept unlearning. Instead of modifying the model across all denoising steps, KSCU selectively applies unlearning to specific denoising steps that contribute most significantly to the target concept’s generation. This targeted approach reduces the number of training iterations required for unlearning, thereby minimizing unnecessary parameter adjustments and preserving the model’s overall generative quality.
3 Key Step Concept Unlearning
As shown in Figure 2, our framework consists of three main components: the Key Step Unlearning Loss module, including the unlearn model that undergoes fine-tuning and the guidance model that provides reverse guidance; the Key Step Table module, which controls the denoising steps during training; and the Prompt Augmentation module to help the model learn more generalized semantic distribution of the unlearned concept in the text embedding space.
3.1 Background
Diffusion models [17] are a class of generative models that learn data distributions by gradually denoising a normally distributed variable. The process consists of two phases: the forward diffusion process and the reverse denoising process. In the forward process, Gaussian noise is incrementally added to the data over a series of time steps, transforming the data into a nearly isotropic Gaussian distribution. Conversely, the reverse process learns to iteratively denoise the data, recovering the original distribution.
Latent Diffusion Models. While traditional diffusion models operate directly in the high-dimensional data space, latent diffusion models (LDMs) [40] introduce a more efficient approach by performing the diffusion process in a compressed latent space. LDMs first employ an autoencoder to encode the input data $\mathbf{x}$ into a lower-dimensional latent representation $\mathbf{z}=E(\mathbf{x})$, where $E$ is the encoder. The diffusion process is then applied to $\mathbf{z}$ instead of $\mathbf{x}$, significantly reducing computational complexity.
The key distinction of LDMs lies in their separation of the diffusion process from the high-dimensional data space. The forward process in the latent space is defined as:
$q(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\mathbf{z}_{t};\sqrt{1-\beta_{t}}\mathbf{z}_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\mathbf{z}_{t}$ is the latent representation at time step $t$. The reverse process learns to denoise the latent representation:
$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},t),\Sigma_{\theta}(\mathbf{z}_{t},t)).$ (2)
After the reverse process, the decoder $D$ maps the denoised latent representation $\mathbf{z}_{0}$ back to the data space: $\mathbf{x}=D(\mathbf{z}_{0})$. This two-stage approach not only reduces computational costs but also enables the generation of high-resolution samples by focusing on a compact latent space.
Classifier-Free Guidance. Classifier-Free Guidance [16] is a technique designed to improve the controllability of diffusion models without relying on an external classifier. Instead of conditioning the model on explicit labels, classifier-free guidance jointly trains a conditional and an unconditional diffusion model. Formally, let $\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ denote the conditional prediction and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ denote the unconditional prediction, where $\mathbf{c}$ is the conditioning information (e.g., a class label or text embedding). The guided prediction is computed as:
$\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))$ (3)
where $w$ is the guidance scale controlling the strength of conditioning. A higher $w$ emphasizes the conditioning information, while $w=0$ reduces the model to an unconditional generator.
3.2 Key Step Unlearning Loss
Figure 3: Sampling with ”nudity woman” prompt reveals emergence of nudity concepts in the final 10 denoising steps.
According to formula (3), in a classifier-free guidance diffusion model, the linear extrapolation of $\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ serves as the gradient of the implicit classifier $\nabla_{x_{t}}\log p(y|x_{t})$. Our objective is not to generate images corresponding to the text prompt $c$, but rather to shift the model’s predicted noise $\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ in the opposite direction of $\nabla_{x_{t}}\log p(y|x_{t})$. Furthermore, we exclusively fine-tune the conditional noise $\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ to ensure that the generation of unrelated concepts remains unaffected. The fine-tuning objective is formulated as:
$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)-w\cdot(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))\end{split}$ (4)
From this, we derive the unlearning loss:
$\begin{split}\text{Loss}_{\text{unlearn}}&=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c})-(\epsilon_{\theta}(\mathbf{x}_{t},t)-\left(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t)\right))\right\|_{2}\\ \end{split}$ (5)
where $\epsilon^{*}_{\theta}$ represents the unlearned model, $\epsilon_{\theta}$ is the guidance model, $x_{t}$ denotes the denoised image at step $t$, and $c$ corresponds to the text prompt.
More importantly, the Key Step Unlearning Loss also accounts for the optimization of unconditional noise. To preserve the model’s ability to generate unrelated concepts, we introduce a regularization term based on unconditional noise prediction to minimize parameter changes. Specifically, we constrain the L2 distance between the unconditional noise predictions of the guidance and unlearned models, offering partial protection for shared parameters. Additionally, we observe that unconditional noise, which uses the previous sampling step as a prior, gradually aligns with the target concept as denoising progresses. As shown in Figure 3, we use the prompt ”nudity woman” with DDIM steps set to 50 and sample images using Stable Diffusion v1-4. From left to right, the images correspond to step = 40, step = 50, and the last 10 steps sampled using only unconditional noise. The qualitative results show that even unconditional noise alone causes the generation to drift toward the ”nudity” concept. This highlights the necessity of applying unlearning not only to the conditional branch but also to the unconditional branch in classifier-free guidance diffusion models. Accordingly, we apply unlearning training to the unconditional component as well. The regularization loss is formulated as:
$\text{Loss}_{\text{regular}}=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t)-((1+\lambda_{1})\epsilon_{\theta}(\mathbf{x}_{t},t)-\lambda_{1}\ \epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c}))\right\|^{2}_{2}$ (6)
where $\lambda_{1}$ needs to be a small value that is linearly related to $t$. For $t$ from 1000 to 0, we set $\lambda_{1}=(1000-t)/(2*10^{5})$
The final Key Step Unlearning Loss function for KSCU is defined as:
$\text{Loss}=\text{Loss}_{\text{unlearn}}+\lambda_{2}\cdot\text{Loss}_{\text{regular}}$ (7)
where $\lambda_{2}=1\times 10^{-4}$ in KSCU.
Algorithm 1 Key Step Table Generation
Input:start step $S$, end step $E$, table length $L$, full loops before shift $loop_{n}$
[…]
1:  $KeyStepTable\leftarrow None$
2:  $s_{cur}\leftarrow S$
3:  $loop_{cur}\leftarrow 0$
4:  while $len(KeyStepTable)<L$ do
5:     $KeyStepTable\leftarrow KeyStepTable\cup[s_{cur},s_{cur}+1,\dots,E]$
6:     $loop_{cur}\leftarrow loop_{cur}+1$
7:     if $loop_{cur}=loop_{n}$ then
8:        $s_{cur}\leftarrow\min(s_{cur}+1,E-1)$
9:        $loop_{cur}\leftarrow 0$
10:     end if
11:  end while
3.3 Key Step Table Generation
The goal of concept unlearning is to prevent the generation of a target concept. Due to the nature of diffusion-based generation, high-frequency details in the image are primarily refined during the later denoising steps. Therefore, we hypothesize that these later steps play a more critical role in the unlearning process. To validate this, we design a preliminary experiment using ESD [12] to fine-tune Stable Diffusion v1-4, while restricting the range of training steps. The results show that, for NSFW content unlearning, training only the last 70% of denoising steps (with the total number of training iterations also reduced to 70% of the original) preserves approximately 95% of the unlearning accuracy observed when training across all steps. This indicates that, compared to early steps, the later steps contribute more significantly to concept forgetting. Limiting unlearning to these steps—those responsible for high-frequency detail generation—can still achieve effective unlearning performance.
Based on the above observation, we design a step selection algorithm to generate Key Step Table. The Key Step Table consists of incrementally growing step intervals and shifts the starting point forward after a fixed number of loops to progressively cover a broader range of late denoising steps. As shown in Algorithm 1, the algorithm constructs the Key Step Table from a given starting step $S$ to an ending step $E$. In each iteration, the full interval from $s_{cur}$ to $E$ is appended to the Key Step Table. After $loop_{n}$ iterations, the starting step $s_{cur}$ is shifted forward by one, and the process is repeated until the Key Step Table reaches the desired length $L$. This mechanism ensures that training focuses more on later denoising steps, which are more influential, while avoiding overfitting to a single segment.
Furthermore, our analysis indicates that the most influential steps vary by concept type. Specifically, for NSFW content and class unlearning, the last 70% of denoising steps are most critical, whereas for style unlearning, only the last 50% of steps play a dominant role, while for individual instance forgetting, this interval can be further reduced to the last 20%. These observations highlight the need for a structured step selection approach.
3.4 Prompt Augmentation
Figure 4: Comparison of Semantic Range in Prompt Embedding Space. Left: Limited effective unlearning semantic range (red dashed boundary) with single target concept prompts. Right: Expanded effective unlearning semantic range (green dashed boundary) using augmented prompts.
For prompt-based concept unlearning in T2I DMs, the objective is to erase undesired visual concepts from the model’s generated results. Prompts describe these visual concepts to explicitly specify the target concept for unlearning. However, accurately depicting a complex visual concept through textual prompts is often challenging. Inaccurate prompts may lead to incomplete unlearning, where related concepts remain, or incorrect removal of unrelated concepts. Recently, many methods [54, 5] have recognized this issue and introduced improvements targeting prompt section. In this section, we propose a simpler yet effective Prompt Augmentation method to address this problem.
For a given concept $c$ to be unlearned, we use large language models such as ChatGPT [32] and DeepSeek [8] to generate ten augmentation rules that expand the original concept into diverse textual descriptions. For example, given the concept ”cat” and the rule ”a picture of {}, we obtain: ”a picture of a cat”. We then introduce perturbations by randomly inserting scrambled characters, shuffling word order, or removing non-keywords. Additionally, Gaussian noise is injected into the text embeddings of these modified descriptions [19]. Since the embeddings are ultimately perturbed, the quality of the augmentation rules need not be strictly controlled.
As shown in Figure 4, in the embedding space [36], nearby vectors share similar semantics and induce similar generative behavior. Leveraging this property, Prompt Augmentation enables the model to better perceive the semantic space of the target concept during unlearning, leading to more effective concept erasure. Theoretically, Prompt Augmentation constructs numerous embeddings close to the target concept, aligning with its semantic distribution. This allows the model to unlearn not only the concept itself but also its semantic variations. In contrast, adversarial samples are discrete and harder to construct. Compared to adversarial training, which is costly in time and computation, Prompt Augmentation offers a more efficient alternative for enhancing unlearning robustness.
4 Experimental Results
Table 1: Class and Style Unlearning Performance on Unlearn Canvas. The bold font represents the best, and the underlined font represents the second-best.
Method class style
[…]
ESD [12] 68.6 98.9 96.4 26.1 100.0 85.7 99.0 25.3
CA [23] 32.5 86.7 83.4 51.9 96.0 86.6 88.7 37.4
FMN [56] 25.9 93.0 96.1 36.8 41.5 95.1 91.3 28.5
EDiff [53] 78.0 94.1 90.1 48.2 67.5 93.7 98.3 39.8
Salun [11] 51.3 93.5 94.6 47.9 17.5 96.3 95.2 34.8
SHS [52] 62.5 54.6 56.7 79.4 65.0 51.1 29.4 65.9
KSCU 70.0 99.0 97.1 25.8 100.0 88.7 99.3 23.1
4.1 Experimental Setup
Dataset. To rigorously evaluate various concept unlearning methods, we conducted experiments on four concept unlearning task types: Class, Style, NSFW, and Instance on two datasets, including the I2P [44] and Unlearn Canvas [57].
I2P is a dataset containing 4,703 NSFW prompts. For our experiments, we selected 931 prompts labeled as ”sexual” to generate images. We then used Nudenet [34] (threshold = 0) to analyze the generated images, detecting 575 nudity regions in the outputs from these 931 prompts. Across the entire I2P dataset, Nudenet identified a total of 793 nudity regions. This dataset setting reduced the computational cost while still preserving the majority of the experimental requirements.
Unlearn Canvas is a recently released dataset comprising 20 classes and 60 distinct styles of high-resolution stylized images. Since testing the entire dataset was computationally intensive, we selected 10 classes and 10 styles for the evaluation, including the following:
  • Class: Architectures, Bears, Cats, Dogs, Humans, Towers, Flowers, Sandwiches, Trees, Sea.
  • Styles: Abstractionism, Artist Sketch, Cartoon, Impressionism, Monet, Pastel, Pencil Drawing, Picasso, Sketch, Van Gogh.
Such the selection also ensured diversity: the chosen classes include animals, plants, natural scenes, and man-made objects, while the styles cover both artistic styles such as ”Monet” and illustrative styles such as ”Sketch”.
Evaluation. In our experiments, we evaluated the unlearning effectiveness of various methods using Unlearn Accuracy (UA) originally from [12]. To achieve a more concise and effective evaluation of unlearning effectiveness, we computed UA as follows: $UA=1-x/y$. In the I2P dataset, $x$ represented the number of nudity regions detected in the output images by the unlearned model, and $y$ was 575. Compared to previous methods [12] that evaluated each exposed region individually, we aggregated all cases into a single metric, Unlearn Accuracy, which offered a more concise, analyzable and convincing measure of unlearning performance. For the Unlearn Canvas dataset, $x$ denoted the number of target concepts detected by the classifier [10], while $y$ corresponded to the total number of generated images based on the target prompts.
We also utilized the Fréchet Inception Distance (FID) [15], a widely used metric for assessing the quality of generated images, which measured the distance between the generated and real images in feature space. In the I2P dataset, we generated images from corresponding prompts using the COCO [24] dataset and computed the FID by comparing them with real images from coco30k to assess how well each baseline preserved the model’s generative capacity. For Unlearn Canvas, we generated images using the prompt ”A {class} image in {theme} style.”, then calculated the FID by comparing these images with real images from the Unlearn Canvas dataset. Besides UA and FID, Unlearn Canvas introduced two additional metrics: In-domain Retain Accuracy (IRA) and Cross-domain Retain Accuracy (CRA), which helped evaluate the model’s ability to retain relevant information. Lastly, we assessed the model’s defense capabilities by constructing adversarial text prompts using the P4D [6] and Unlearndiff(UDA) [58] methods based on I2P prompts, comparing them using Attack Success Rate (ASR).
Figure 5: The figure demonstrates KSCU’s unlearning effectiveness and generative retainability on Unlearn Canvas. Images within the red border show the removal of target concepts, while images outside the red border highlight KSCU’s generative retainability. KSCU only aims to remove the target concept without enforcing constraints on the final output. In Appendix, we provide a controllable version of KSCU that can replace the target concept with any specified alternative concept.
Baseline and Training Details. To provide a comprehensive comparison of various methods, we selected six state-of-the-art (SOTA) unlearning techniques, including ESD [12], CA [23], FMN [56], EDiff [53], Salun [11], and SHS [52]. For KSCU’s training setup, we used a batch size of 1. The scheduler was set to 50 timesteps. For style unlearning tasks, the Key Step Table was configured to start from step 25 and end at step 50. For other tasks, the training Key Step Table began at step 15. We employed the Adam optimizer with a learning rate of $1\times 10^{-5}$, applying a total of 750 parameter updates. For all baseline methods, we adopted the default hyperparameters as specified in their original papers. We employed DDIM for denoising, using 50 steps. For Unlearn Canvas, we used Stable Diffusion 1.5, while for all other experiments, unless explicitly stated otherwise, Stable Diffusion 1.4 was used. All experiments were conducted on a single Nvidia L40 GPU.
4.2 Class and Style Unlearning
To evaluate the effectiveness of our proposed KSCU, we first performed qualitative experiments on class and style unlearning tasks. As shown in Figure 5, for different image samples, KSCU has completely blocked the generation of the target class or style, with minimal impact on unrelated classes and styles. Then we quantitatively evaluated its performance, as described below.
Class Unlearning. In the selection of test categories for class unlearning, we randomly selected ten classes encompassing a diverse range from Unlearn Canvas, including animals, plants, man-made objects, and natural elements. The experimental results presented in Table 1 indicated that KSCU achieved superior category-wide unlearning while better preserving the model’s generative capability—an advantage not observed in other SOTAs.
Specifically, as illustrated in Table 1, KSCU achieved 70%, 99.0%, 97.1%, and 25.8 on UA, IRA, CRA, and FID, respectively. In terms of UA, KSCU ranked as the second-best method, while it achieved the best performance on IRA, CRA, and FID. A method that aggressively removed target concepts might achieve higher UA but often at the cost of severely compromising image quality and diversity. Conversely, an overly conservative method may ensure generative retainability but fail to achieve effective unlearning. Therefore, evaluating a T2I diffusion model unlearning method requires consideration of both unlearning effectiveness (UA) and generative retainability (IRA, CRA, FID). In these two regards, KSCU demonstrates a more favorable trade-off. A high UA of 70% indicates strong unlearning effectiveness, while the best IRA and CRA scores confirm its superior ability to preserve generative diversity. Furthermore, KSCU outperforms all other SOTAs in FID, indicating that its unlearned model can generate higher-quality images, further validating its generative retainability. Conversely, while EDiff attained the highest UA, its FID reached 48, suggesting that although it exhibited strong unlearning performance, it significantly degraded the model’s generative ability. This low generative retainability of EDiff indicates that it is difficult to use in practical applications.
Style Unlearning. We conducted experiments on image style unlearning using the Unlearn Canvas dataset. We quantitatively assessed the feasibility of removing artistic styles to address copyright concerns using the IRA and CRA metrics. Given computational constraints, we selected 10 styles from Section 4.1, covering three artistic styles and seven artistic representations. This selection ensured a diverse evaluation of style unlearning across different visual attributes while maintaining a feasible training cost. The experimental results presented in Table 1 indicated that KSCU achieved strong unlearning effectiveness and retainability, while maintaining a favorable balance between the two on the style unlearning task.
Specifically, as shown in Table 1, KSCU achieved scores of 100%, 88.7%, 99.3%, and 23.1 in UA, IRA, CRA, and FID, respectively. Among these, UA, CRA, and FID were the best across all compared methods. While KSCU’s IRA was slightly lower than the Salun method with very low unlearning effectiveness (UA $=$ 17.5%), this method reached high IRA at the expense of insufficient concept forgetting. In contrast, when compared with other high unlearning effectiveness methods such as CA and ESD (UA $>$ 95%), KSCU demonstrated clearly superior IRA, CRA, and FID. These results indicated that KSCU not only effectively eliminated traces of artistic styles but also better preserved the model’s ability to generate unrelated concepts.
KSCU halved the iteration count compared to ESD by selectively fine-tuning only the latter 50% of denoising steps. Despite reducing the number of parameter updates by half,, KSCU still outperformed ESD in style unlearning, reinforcing that full-step fine-tuning is unnecessary for concept unlearning in T2I diffusion models. Focusing on the most influential steps enables more efficient and controlled unlearning while minimizing disruptions to the model’s generative capacity. These findings suggest that the Key Step Table optimization strategy adopted by KSCU offers a more practical and scalable solution for concept unlearning in large-scale diffusion models.
Table 2: Evaluation of Unlearning Capability on ”Nudity”. The bold font represents the best, and the underlined font represents the second-best.
Method UA(%$\uparrow$ P4D(%$\downarrow$ UDA(%$\downarrow$ FID-30$\downarrow$
ESD [12] 88.2 39.3 73.7 15.4
CA [23] 34.3 - - 15.3
Salun [11] 100.0 0.0 4.2 67.9
FMN [56] 38.6 - - 20.9
EDiff [53] 93.9 29.6 52.5 19.8
SHS [52] 84.3 62.2 95.5 18.1
KSCU 96.5 23.2 47.2 14.1
4.3 NSFW Unlearning
NSFW Unlearning. In the NSFW concept forgetting experiment, we fine-tuned the models using ”nudity” as the target concept and obtained the corresponding unlearned models. Each method used its corresponding unlearned model to generate 931 and 30,000 images based on prompts from the i2p [44] and COCO [24] datasets, respectively. Table 2 presents the results, assessing the unlearning effectiveness of various methods using Unlearn Accuracy, P4D, and UDA’s ASR. FID is also compared to evaluate generative retainability. Note that CA and FMN, which fine-tune only the cross-attention module, show limited NSFW unlearning effects and are not further tested for robustness against attacks.
As shown in Table 2, KSCU balanced unlearning effectiveness and generative retainability, achieving a UA of 96.5%, surpassing other methods in forgetting capability. Meanwhile, its FID-30K of 14.1 was the lowest among all baselines, confirming that KSCU best generative retainability. In contrast, although Salun achieves the highest UA, its FID of 67.9 reveals a significant degradation in generative quality. Moreover, we observed extensive large-area distortions in Salun’s generated results, indicating severe over-unlearning. Compared to methods with similar FID-30K scores, such as ESD and CA, KSCU demonstrated superior forgetting performance. Conversely, while methods like EDiff and Salun matched KSCU in unlearning effectiveness, they fell short in retaining generative ability. These results established KSCU as the most effective and balanced approach for NSFW concept unlearning.
Robustness. In the adversarial sample defense experiment, excluding Salun (which exhibited over-unlearning), KSCU achieved the lowest ASR against P4d [6] and UDA [58] among all other SOTAs, demonstrating its strong robustness. On the other hand, we observed a certain degree of correlation between the performance of different methods under adversarial attacks and their forgetting precision. We believed this was because the i2p prompts used in the experiment were manually selected, diverse prompts with inherent adversarial properties. As a result, the unlearning performance on i2p can, to some extent, reflect the robustness of a given method in complex real-world scenarios or against adversarial samples.
4.4 Instance Unlearning
Figure 6: KSCU vs. ESD in Instance Unlearning. KSCU achieves forgetting with only 20% of ESD’s iterations. Despite ESD’s greater adjustments, KSCU demonstrates superior performance in erasing instance information.
To demonstrate KSCU’s effectiveness in unlearning instances memorized by diffusion models, we qualitatively compared the unlearning performance of KSCU and ESD through visualized instance unlearning results. The reason for choosing ESD was that the architecture of ESD was more similar to ours and ESD had shown good performance in previous experiments.
As shown in Figure 6, KSCU demonstrated superior performance in unlearning public figures. In KSCU’s results, facial features of public figures were effectively altered, whereas ESD partially retained the original characteristics. For artificial objects, ESD’s unlearning effect was barely noticeable, whereas KSCU replaced the target with either a visually similar but semantically distinct object (e.g., Eiffel Tower $\rightarrow$ highway) or a semantically related alternative (e.g., Titanic $\rightarrow$ small boat). KSCU, due to its well-designed approach, achieved a more effective instance unlearning performance. More importantly, KSCU achieved this with lower computational cost, highlighting its potential for broader applications.
4.5 Ablation Study
Table 3: Ablation Study: Impact of Prompt Augmentation(PA) and Key Step Table (with Different Start $S$) on KSCU’s Performance
Method UA(%$\uparrow$ P4D(%$\downarrow$ UDA(%$\downarrow$ FID-30$\downarrow$
ESD [12] 88.2 39.3 73.7 15.4
KSCU w/o PA 92.5 30.3 51.8 14.5
KSCU $S$=20 79.7 59.2 85.3 15.7
KSCU $S$=10 96.1 23.0 48.1 14.3
KSCU $S$=0 96.0 26.7 54.5 18.8
KSCU 96.5 23.2 47.2 14.1
In the ablation study, we mainly evaluated the impact of the Key Step Table and Prompt Augmentation. After removing Prompt Augmentation, KSCU exhibited a decrease in UA and a significant increase in ASR against adversarial examples, indicating a reduction in model robustness. This strongly confirmed that Prompt Augmentation enhanced the model’s ability to better perceive the semantic space of the target concept, achieve precise concept erasure, and improve robustness.
In the ablation experiment of the Key Step Table’s start, we observed that when $S$=10 (covering 80% of steps), the performance was similar to $S$=15 (covering 70% of steps). However, when $S$=20 (covering 60% of steps), the unlearning effect of KSCU significantly decreased, suggesting that steps beyond 15 played a key role in the imaging of “nudity.” In the $S$=0 (covering 100% of steps) experiment, the unlearning performance remained strong. However, since all steps were affected, the FID increased, indicating a more significant negative impact on the model’s generative ability. Notably, the unlearning effectiveness of $S$=0 (100% steps), $S$=10 (80% steps), and $S$=15 (KSCU) was highly similar. This suggested that the first 30% of steps had little influence on the final rendering of ”nudity,” making their unlearning unnecessary. The results of the ablation experiment further demonstrated that for concept unlearning, KSCU, which focuses on unlearning key steps, was superior to the previous method of adjusting parameters for all steps.
5 Conclusion
This study investigates the impact of denoising steps on concept unlearning in diffusion models and proposes KSCU, a method that targets key denoising steps for concept unlearning. Through experiments on four different concept unlearning tasks, we demonstrate that KSCU surpasses previous state-of-the-art methods in terms of unlearning effectiveness, robustness, and generative retainability. The core ideas behind the key steps introduced in this work provide valuable insights for research in related fields.
Appendix
Appendix A Analysis of Frequency Component Destruction in Diffusion Models
In score-matching diffusion models, data undergoes a forward diffusion process, where noise is progressively added until only pure noise remains, followed by a reverse denoising process, where the model learns to reconstruct the original data. Since the reverse process restores information corresponding to the degradation in the forward process, proving that diffusion models first sample low-frequency components and later refine high-frequency details requires showing that high-frequency components are the first to be destroyed in the forward process.
A.1 Definition of the Diffusion Process
We consider a continuous-time diffusion process, where the initial data $\mathbf{x}_{0}$ evolves into $\mathbf{x}_{t}$ over time $t\in[0,T]$. The diffusion process is governed by the following stochastic differential equation (SDE):
$d\mathbf{x}_{t}=f(\mathbf{x}_{t},t)dt+g(t)d\mathbf{w}_{t}$ (8)
where: - $\mathbf{w}_{t}$ represents Brownian motion, modeling random noise. - $f(\mathbf{x}_{t},t)$ is the drift term, which we set to zero for simplification. - $g(t)$ is the time-dependent diffusion coefficient, controlling noise intensity.
Under these assumptions, the simplified diffusion process is given by:
$d\mathbf{x}_{t}=g(t)d\mathbf{w}_{t},\quad\mathbf{x}_{0}\text{ as the initial condition.}$. (9)
A.2 Frequency-Domain Analysis
To analyze how different frequency components evolve during diffusion, we apply the Fourier transform to convert the signal from the time domain to the frequency domain. Let $\hat{\mathbf{x}}_{t}(\omega)$ denote the Fourier transform of $\mathbf{x}_{t}$, where $\omega$ represents frequency. The linearity of the Fourier transform and the properties of the Wiener process enable a more straightforward analysis in the spectral domain.
Since $d\mathbf{w}_{t}$ represents white noise, its power spectral density is constant across frequencies.
A.3 Derivation of Frequency Component Degradation
A.3.1 Representation of the Initial Signal and Noise
During the forward diffusion process, the signal $\mathbf{x}_{t}$ can be expressed as the sum of the initial signal $\mathbf{x}_{0}$ and accumulated noise:
[…]
In the frequency domain, the Fourier transform of $\mathbf{x}_{t}$ is given by:
$\hat{\mathbf{x}}_{t}(\omega)=\hat{\mathbf{x}}_{0}(\omega)+\int_{0}^{t}g(s)\hat{d\mathbf{w}}_{s}(\omega).$ (11)
Here, $\hat{d\mathbf{w}}_{s}(\omega)$ is the Fourier transform of the noise term, satisfying: - $\mathbb{E}[\hat{d\mathbf{w}}_{s}(\omega)]=0$, - $\mathbb{E}[|\hat{d\mathbf{w}}_{s}(\omega)|^{2}]=ds$.
Since $d\mathbf{w}_{t}$ is white noise, its power spectral density remains uniform across all frequencies.
A.3.2 Power Spectral Density of the Signal
The power spectral density of $\mathbf{x}_{t}$ is given by:
$\mathbb{E}[|\hat{\mathbf{x}}_{t}(\omega)|^{2}]=\mathbb{E}[|\hat{\mathbf{x}}_{0}(\omega)+\int_{0}^{t}g(s)\hat{d\mathbf{w}}_{s}(\omega)|^{2}].$ (12)
Expanding the expectation:
$\begin{split}\mathbb{E}[|\hat{\mathbf{x}}_{t}(\omega)|^{2}]=&|\hat{\mathbf{x}}_{0}(\omega)|^{2}+\mathbb{E}\left[\left|\int_{0}^{t}g(s)\hat{d\mathbf{w}}_{s}(\omega)\right|^{2}\right]\\ &+2\text{Re}\left\{\hat{\mathbf{x}}_{0}(\omega)\mathbb{E}\left[\int_{0}^{t}g(s)\hat{d\mathbf{w}}_{s}(\omega)^{*}\right]\right\}.\end{split}$ (13)
Since $\mathbb{E}[\hat{d\mathbf{w}}_{s}(\omega)]=0$, the cross-term vanishes, leaving:
[…]
A.3.3 Signal-to-Noise Ratio (SNR) Analysis
The signal-to-noise ratio (SNR) is defined as:
[…]
Since noise power is independent of frequency, SNR evolution is dictated entirely by the initial power spectral density $|\hat{\mathbf{x}}_{0}(\omega)|^{2}$.
A.3.4 Frequency Dependence of SNR Decay
For natural signals (e.g., images and audio), the power spectrum typically follows a power-law decay:
$|\hat{\mathbf{x}}_{0}(\omega)|^{2}\propto\frac{1}{\omega^{\alpha}},\quad(\alpha>0).$ (16)
As $\omega$ increases, $|\hat{\mathbf{x}}_{0}(\omega)|^{2}$ rapidly decreases, implying that high-frequency components exhibit lower SNR.
[…]
Since $|\hat{\mathbf{x}}_{0}(\omega)|^{2}$ decreases with $\omega$, $t_{\text{th}}$ is smaller for higher frequencies, meaning high-frequency information is lost earlier.
A.4 Conclusion
The analysis confirms that, in the forward diffusion process, high-frequency components are degraded first, followed by lower-frequency components. Consequently, in the reverse denoising process, diffusion models reconstruct low-frequency structures first before refining high-frequency details. This frequency-dependent degradation provides theoretical justification for the progressive nature of generative sampling in diffusion models.
Appendix B More Experimental Details
The models used in our experiments primarily include Stable Diffusion v1-5 and Stable Diffusion v1-4. KSCU applies targeted fine-tuning based on different unlearning tasks:
  • Class unlearning: Fine-tuning is conducted on the last 70% of denoising steps with 700 iterations and a batch size of 1.
  • Style unlearning: Fine-tuning is applied to the last 50% of denoising steps with 500 iterations and a batch size of 1.
  • Instance unlearning: Only the last 20% of denoising steps are fine-tuned, using 200 iterations.
  • NSFW (nudity) unlearning: Fine-tuning is applied to the last 70% of denoising steps with 750 iterations. Unlike other tasks, this process updates all model components except the cross-attention module.
All class, style, and instance unlearning tasks involve fine-tuning the model’s cross-attention layers, whereas the NSFW unlearning task extends fine-tuning to all components except cross-attention.
For experiments on Unlearn Canvas, we use the official implementation provided by Unlearn Canvas for all methods except KSCU and ESD. To simplify the experiments while ensuring their validity, we selected 10 classes and 10 styles for evaluation. We carefully ensured diversity and representativeness in the chosen classes and styles. For experiments on Unlearn Canvas, we followed the original paper and used Stable Diffusion version 1.5. However, unlike the original approach, we computed FID using images generated by the frozen model instead of real data. For class and style unlearning, we train 10 models per method, each generating images based on predefined prompts, resulting in a total of $10\times 51\times 20=10,200$ images. When computing FID, we exclude images corresponding to the 10 target categories entirely and compare the remaining images with those generated by a frozen model.
For experiments on I2P, we employed Stable Diffusion version 1.4. The primary reason for this choice is that, compared to later versions, SDv1.4 generates more ”nudity” images from I2P prompts, making it more suitable for evaluating the effectiveness of different unlearning methods.
For instance unlearning experiments, we conducted our study using Stable Diffusion version 1.4. We present four comparative cases: two involving human subjects and two involving objects. Although we did not perform a quantitative comparison, qualitative results indicate that KSCU outperforms ESD. For instance unlearning tasks, modifying high-frequency information alone is sufficient to transform an instance into another. The design of KSCU makes it particularly well-suited for such tasks.
The visualization results for Stable Diffusion version 1.5 and later versions are presented in the subsequent sections.
Appendix C Target Concept Unlearning
In Section 3.2, we derived our loss function, which ensures that the model samples in the direction opposite to the original concept $c$. This approach effectively prevents the generation of $c$ by replacing it with an unrelated concept. However, the replacement concept is uncontrolled, leading to unpredictable visual outputs. To address this, we introduce a slight modification to the loss function, guiding the model to replace $c$ with a specified target concept $c^{-}$, rather than an arbitrary alternative. The modified formulation is as follows:
$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{-}}\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{+}}\end{split}$ (18)
where $c^{+}$ represents the concept to be unlearned, and $c^{-}$ denotes the target replacement concept. By applying the following modified loss function, we explicitly enforce the substitution of $c^{+}$ with $c^{-}$:
$\text{Loss}_{\text{target unlearn}}=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})\right\|.$ (19)
This formulation allows precise control over the replacement concept, ensuring that the model not only forgets $c^{+}$ but also consistently generates the desired substitute $c^{-}$, thereby enabling targeted concept substitution in diffusion models.
In the Figure 8, we present the results of KSCU’s target concept unlearning.
Appendix D Multi-concept Unlearning
To evaluate the effectiveness of KSCU in multi-concept unlearning, we provide visualized results 7. Our observations indicate that KSCU maintains strong performance in removing multiple concepts.
Figure 7: Results of multi-concept unlearn.
Appendix E More Visualization Results
Figure 8: Visualization results of KSCU with target replacement. The left side is SDv1.5, and the right side is the result of KSCU. The target concepts are bird and Van Gogh, and the replacement concepts are balloon and sketch.
Figure 9: KSCU’s visualization results for “nudity”, with SDv2.1 on the left and KSCU on the right.
Figure 10: visualization results of KSCU about artist styles, SDv2.1 on the left and KSCU on the right.
Figure 11: Visualization results of artist style erasing, model is SDv1.5

### v3

@@
 Abstract
-Text-to-image diffusion models (T2I DMs), represented by Stable Diffusion, which generate highly realistic images based on textual input, have been widely used.
-However, their misuse poses serious security risks.
-While existing concept unlearning methods aim to mitigate these risks, they struggle to balance unlearning effectiveness with generative retainability.To overcome this limitation, we innovatively propose the Key Step Concept Unlearning (KSCU) method, which ingeniously capitalizes on the unique stepwise sampling characteristic inherent in diffusion models during the image generation process.
-Unlike conventional approaches that treat all denoising steps equally, KSCU strategically focuses on pivotal steps with the most influence over the final outcome by dividing key steps for different concept unlearning tasks and fine-tuning the model only at those steps.
-This targeted approach reduces the number of parameter updates needed for effective unlearning, while maximizing the retention of the model’s generative capabilities.Through extensive benchmark experiments, we demonstrate that KSCU effectively prevents T2I DMs from generating undesirable images while better retaining the model’s generative capabilities.
-Our code will be released.
-Preprint. Under review.
+Text-to-image diffusion models (T2I DMs), represented by Stable Diffusion, which generate highly realistic images based on textual input, have been widely used, but their flexibility also makes them prone to misuse for producing harmful or unsafe content.
+Concept unlearning has been used to prevent text-to-image diffusion models from being misused to generate undesirable visual content.
+However, existing methods struggle to trade off unlearning effectiveness with the preservation of generation quality.
+To address this limitation, we propose Key Step Concept Unlearning (KSCU), which selectively fine-tunes the model at key steps to the target concept.
+KSCU is inspired by the fact that different diffusion denoising steps contribute unequally to the final generation.
+Compared to previous approaches, which treat all denoising steps uniformly, KSCU avoids over-optimization of unnecessary steps for higher effectiveness and reduces the number of parameter updates for higher efficiency.
+For example, on the I2P dataset, KSCU outperforms ESD by 8.3% in nudity unlearning accuracy while improving FID by 8.4%, and achieves a high overall score of 0.92, substantially surpassing all other SOTA methods.
+I Introduction
+In recent years, the advancement of text-to-image (T2I) diffusion models [1, 2, 3, 4, 5, 6] has led to their widespread adoption across various domains [7, 8, 9, 10, 11], including short videos, comics, and illustrations.
+These models enable users to translate natural language prompts into high-quality, semantically coherent visual outputs, greatly lowering the entry barrier for creative applications.
+As a result, AI-generated images have become an integral part of daily life for many individuals.
+However, large-scale T2I models are trained on extensive datasets [12, 13, 14, 7] that inevitably contain sensitive or problematic content, such as unauthorized artistic works [15, 16], culturally biased representations, or Not Safe For Work (NSFW) material.
+Consequently, these models can inadvertently regenerate infringing or inappropriate content when prompted with relevant text inputs [17].
+This capability raises pressing concerns regarding copyright infringement, ethical use, and social security risks, highlighting the urgent need for effective mitigation strategies.
 Figure 1: Compared to previous methods (b), KSCU (c) focuses exclusively on the denoising steps that have the most significant impact on concept generation and achieves effective forgetting of the target concept while better preserving the generation of unrelated concepts.
-1 Introduction
-In recent years, the advancement of text-to-image (T2I) diffusion models [22, 20, 9, 16, 26, 43] has led to their widespread adoption across various domains [38, 31, 42, 1, 27], including short videos, comics, and illustrations.
-As a result, AI-generated images have become an integral part of daily life for many individuals.
-However, large-scale T2I models are trained on extensive datasets [45, 4, 37, 38] that often contain sensitive content, such as unauthorized artistic works [46, 47] or Not Safe For Work (NSFW) material.
-Consequently, these models can inadvertently regenerate infringing content when prompted with relevant text inputs [3].
-This raises significant concerns regarding copyright infringement and social security risks, highlighting the urgent need for effective mitigation strategies.
 The generation of NSFW content poses a major security risk for mainstream text-to-image diffusion models.
-Researchers have explored various approaches to mitigate this issue, including dataset-based screening, content filtering [39], and fine-tuning techniques [7, 30, 41].
-concept unlearning, as a fine-tuning-based approach [23, 11, 56, 13, 14, 53, 18], selectively adjusts partial model parameters to remove the model’s ability to generate specific concepts.
-Through concept unlearning, T2I diffusion models effectively ‘forget’ designated visual concepts by replacing them with semantically neutral content.
-This targeted parameter adjustment (referred to as ”concept erase” [12, 21, 14]) provides distinct advantages over conventional methods—offering greater robustness than post-hoc filtering [39] while avoiding the intensive labor of dataset curation [40], and has become one of the most promising approaches to curb the generation of NSFW content in T2I propagation models.
-Although significant progress has been made in concept unlearning for T2I diffusion models, several challenges remain unresolved.
-a.
-Blind unlearning of the entire diffusion process.
-Existing concept unlearning approaches for T2I diffusion models fail to notice for the stepwise denoising characteristics of diffusion models.
-Since diffusion models generate images through a progressive denoising process—where earlier steps primarily capture low-frequency structural information and later steps refine high-frequency details [17, 50, 55, 35, 2]—unlearning does not necessarily require fine-tuning across all denoising steps.
-b.
-Unbalanced unlearning effectiveness and generative retainability.
-Most existing methods require extensive modifications to the original model parameters to ensure unlearning effectiveness.
-However, excessive parameter updates may lead to over-unlearning, negatively affecting the model’s ability to generate unrelated concepts.
-Conversely, strictly limiting the scale of parameter updates to preserve generative retainability—i.e., the model’s ability to retain generation quality for unrelated concepts—makes it difficult to fully unlearn the target concept.
-c.
-Robustness at high cost.
-Adversarial attack [51, 6, 58] research on unlearned models has demonstrated that carefully crafted adversarial text prompts can successfully recover unlearned concepts.
-While adversarial training [21, 59] can enhance robustness against such attacks, it is computationally expensive and brings low generative retainability.
-To address existing challenges, we propose Key Step Concept Unlearning (KSCU), a novel framework for concept unlearning in T2I diffusion models.
-Unlike prior methods that operate across all denoising steps, KSCU selectively fine-tunes a subset of key steps based on a predefined Key Step Table.
-For different unlearning tasks, multiple step intervals are empirically defined, and a selection algorithm prioritizes steps with the greatest influence on the final output, enabling efficient and targeted unlearning.
-To improve robustness, KSCU applies Prompt Augmentation in the final training stage, including random shuffling, word removal, character insertion, and embedding noise.
-These augmented prompts help the model better capture the semantic distribution of the unlearned concept in the embedding space.
-Additionally, we introduce a Key Step Unlearning Loss tailored to classifier-free guidance DMs.
-By combining Key Step Table, Prompt Augmentation, and Key Step Unlearning Loss, KSCU achieves more effective concept unlearning while better ensuring generative retainability.
+To mitigate this issue, researchers have explored various approaches, including dataset-based screening, post-hoc content filtering [18], and fine-tuning techniques [19, 20, 21].
+While dataset curation attempts to remove harmful samples at the source, it is costly and incomplete, as sensitive content is often interwoven with valuable training data.
+Filtering mechanisms at inference time can suppress unsafe outputs but fail to prevent the model from retaining harmful capabilities internally.
+By contrast, concept unlearning has emerged as a promising fine-tuning-based solution [22, 23, 24, 25, 26, 27, 28], which selectively adjusts partial model parameters to remove the model’s ability to generate specific concepts.
+Through concept unlearning, T2I diffusion models effectively ‘forget’ designated visual concepts by replacing them with semantically neutral or degraded alternatives.
+This targeted parameter adjustment (also known as concept erasure [29, 30]) offers distinct advantages over conventional approaches—providing greater robustness than post-hoc filtering [18] while avoiding the intensive labor of dataset curation [31].
+Consequently, concept unlearning has become one of the most promising strategies to curb the generation of NSFW content in diffusion models.
+Although concept unlearning for T2I diffusion models has made notable progress, key challenges remain.
+(a) Blind unlearning across all diffusion steps.
+Existing methods often overlook the stepwise nature of denoising in diffusion models.
+The generation of visual concepts is hierarchical: early steps capture coarse structures and low-frequency semantics, while later steps refine fine-grained textures and high-frequency details [32, 33, 34, 35, 36].
+Ignoring this temporal structure leads to unnecessary parameter updates and reduced efficiency.
+We hypothesize and empirically validate that effective concept unlearning can be achieved by selectively modifying only a small set of key steps that significantly influence the final output.
+(b) Unbalanced unlearning effectiveness and generative retainability.
+Existing approaches often face a trade-off: aggressive parameter updates can achieve stronger forgetting but degrade the generation of unrelated concepts (over-unlearning), while conservative updates preserve retainability but result in incomplete removal of the target concept.
+Achieving a principled balance between these two objectives remains an open challenge.
+To address these challenges, we propose Key Step Concept Unlearning (KSCU), a novel framework for concept unlearning in T2I diffusion models.
+Unlike prior methods that operate across the entire denoising trajectory, KSCU selectively fine-tunes only a subset of key steps determined by a predefined Key Step Table, as illustrated in Figure 1.
+For different unlearning tasks, multiple step intervals are empirically defined, and a step-selection algorithm prioritizes those with the greatest influence on the final output, enabling efficient and targeted unlearning.
+To further enhance robustness, KSCU applies Prompt Augmentation in the final training stage, where augmented prompts help the model better capture the semantic distribution of the unlearned concept in the embedding space.
+Additionally, we introduce a Key Step Unlearning Optimization tailored to Classifier-Free Guidance (CFG) diffusion models, ensuring that the erasure process interacts smoothly with guidance scaling.
+By combining Key Step Table, Prompt Augmentation, and Key Step Unlearning Optimization, KSCU achieves more effective concept unlearning while better preserving generative retainability.
 Our contributions can be summarized as follows:
-  • We propose KSCU, a novel concept unlearning framework for T2I DMs.
-Unlike prior work, KSCU only performs fine-tuning for the key step, achieving efficient unlearning by applying smaller adjustments to the model, thereby maintaining the model’s generation ability, taking both unlearning effectiveness and generative retainability into account.
-  • Our proposed KSCU integrates three innovations: Key Step Table, a step selection strategy focusing on critical denoising steps to achieve effective concept unlearning with minimal fine-tuning overhead; Prompt Augmentation to enhance robustness by introducing diverse prompt variations; and Key Step Unlearning Loss, a concept unlearning loss tailored for classifier-free guided diffusion models, enabling effective concept removal while preserving generation quality.
-  • Extensive experimental results on multiple datasets demonstrate the superiority of KSCU.
-For instance, in the NSFW unlearning experiment, KSCU achieves a remarkable unlearning accuracy of 96.5% with FID of just 14.1, significantly outperforming the previous SOTAs.
-2 Related Work
-2.1 Text-to-Image Diffusion Models
+  • We propose KSCU, a novel concept unlearning framework for T2I diffusion models.
+Unlike prior works, we explicitly observe the hierarchical concept formation process in diffusion and demonstrate that unlearning can be effectively achieved by updating only a small set of key steps.
+KSCU fine-tunes the model exclusively at these steps with minimal adjustments, enabling efficient erasure while preserving generative retainability.
+  • KSCU integrates three key innovations: Key Step Table, a step-selection strategy that focuses on critical denoising steps to achieve effective unlearning with minimal overhead; Prompt Augmentation, which enhances robustness by enriching semantic coverage through diverse prompt variations; and Key Step Unlearning Optimization, an optimization scheme tailored for CFG-based diffusion models, enabling effective concept removal while maintaining high image quality.
+  • We conduct extensive experiments on multiple benchmarks, covering NSFW, style, and category unlearning.
+Results show that KSCU consistently outperforms state-of-the-art baselines.
+For instance, in NSFW unlearning, KSCU achieves an overall score of 0.92 and an unlearning accuracy of 96.5% with an FID of 14.1, significantly surpassing previous SOTAs.
+II Background & Related Work
+II-A Text-to-Image Diffusion Models
 Diffusion models [37] were initially proposed as a generative framework based on the reverse denoising process to reconstruct target input images.
-With the introduction of advanced techniques such as score-based models [50], noise-conditioned score network [49], and denoising-based guidance [9], these models have become increasingly popular for text-to-image (T2I) generation.
-The emergence of latent diffusion models (LDMs) [40] and classifier-free guidance [16] has further improved the controllability and efficiency of diffusion-based T2I systems.
-Their ability to generate high-quality, semantically coherent images has established diffusion models as the dominant approach in T2I generation.
-Building on these advancements, recent research has focused on optimizing T2I diffusion models in terms of efficiency, scalability, and applicability.
+Compared with earlier paradigms such as GANs [38] and VAEs [39], diffusion-based approaches demonstrate superior training stability, more faithful mode coverage, and better scalability to high-dimensional distributions.
+With the introduction of advanced techniques such as score-based models [40], noise-conditioned score networks [41], and denoising-based guidance [3], these models have become increasingly popular for text-to-image (T2I) generation.
+The emergence of Latent Diffusion Models (LDMs) [31] has further improved scalability and efficiency.
+By operating in a learned latent space, LDMs substantially reduce the computational burden while retaining high generative fidelity, making them feasible for large-scale training.
+Specifically, given an image $\mathbf{x}$, an encoder $E(\cdot)$ maps it to the latent code $\mathbf{z}=E(\mathbf{x})$.
+The forward diffusion process then gradually corrupts the latent with Gaussian noise over $T$ steps:
+$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
+where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ is the cumulative product of noise schedule coefficients.
+During inference, the model reverses this process by predicting the added noise $\boldsymbol{\epsilon}$ at each step and updating the latent accordingly:
+$\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\hat{\boldsymbol{\epsilon}}_{\theta}\right)+\sigma_{t}\,\boldsymbol{\eta},\quad\boldsymbol{\eta}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (2)
+where $\hat{\boldsymbol{\epsilon}}_{\theta}$ is the predicted noise, and $\sigma_{t}$ is typically set according to the scheduler.
+To enhance controllability, Classifier-Free Guidance (CFG) [4] interpolates between conditional and unconditional predictions:
+$\hat{\boldsymbol{\epsilon}}_{\theta}=\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t)+w\cdot\left(\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t)\right),$ (3)
+where $w$ denotes the guidance scale, and $\boldsymbol{\epsilon}_{\theta}(\cdot)$ and $\boldsymbol{\epsilon}_{\theta}(\cdot,\mathbf{c})$ represent the unconditional and conditional predictions, respectively.
+Together, LDMs and CFG have established diffusion models as the dominant approach in T2I generation.
+Their ability to generate high-quality, semantically coherent images, even under complex compositional prompts, has set a new benchmark in the field.
+On this basis, recent research has focused on optimizing T2I diffusion models in terms of efficiency, scalability, and applicability.
 Significant breakthroughs have been made in acceleration algorithms [42, 43, 44, 45], model architectures [31], and large-scale datasets [12, 13, 14, 7], all of which have contributed to the widespread adoption of these models in practical applications [7, 8, 9, 10].
+Beyond visual quality, diffusion models have also demonstrated strong generalization in multimodal and cross-domain settings, such as image editing, 3D synthesis, and video generation [46, 47].
 However, despite their advantages, these models are vulnerable to misuse, leading to concerns over NSFW content [18], unauthorized depictions of individuals, and copyright-infringing artwork [10].
+Furthermore, the sheer expressive power of diffusion models introduces unique risks, such as prompt injection attacks, data leakage, and ethical dilemmas in human–AI collaboration.
 As T2I diffusion models continue to advance in realism and fidelity, addressing their ethical and legal challenges has become an urgent priority for the research community.
-Figure 2: The KSCU framework consists of three key modules: Key Step Unlearning Loss, Key Step Table, and Prompt Augmentation.
-The Key Step Table predefines step selection strategies for four concept unlearning tasks (Class, Style, NSFW, and Instance).
+Efforts in controllable generation, safety-aware fine-tuning, and concept erasure are increasingly viewed as essential to ensure their responsible deployment.
+Figure 2: The KSCU framework consists of three key modules: Key Step Table, Prompt Augmentation, and Key Step Unlearning Optimization.
 During training, a denoising step $t$ is selected from the Key Step Table.
-The unlearn model performs denoising sampling to obtain $x_{t}$, while the Prompt Augmentation module converts the target concept $c$ into an augmented prompt $c^{*}$.
-Using $x_{t}$, $t$, $c$, and $c^{*}$, the Key Step Unlearning Loss is computed via Equation (7), and designated parameters of the unlearn model are updated through backpropagation.
-2.2 Diffusion Model Concept Unlearning
+The unlearn model performs denoising sampling to obtain $z_{t}$, while the Prompt Augmentation module converts the target concept $c$ into an augmented prompt $c^{*}$.
+Using $z_{t}$, $t$, $c$, and $c^{*}$, the Key Step Unlearning Optimization is computed via Equation (7), and designated parameters of the unlearn model are updated through backpropagation.
+II-B Diffusion Model Concept Unlearning
 Large-scale datasets [12, 13, 14, 7] often contain undesirable data, including content with inherent biases, ethically problematic information, or copyrighted materials that may later become restricted.
-Models trained on such datasets acquire the ability to generate this sensitive content, posing significant security and ethical risks in text-to-image (T2I) generation.
-To mitigate these concerns, concept unlearning in diffusion models has been proposed, aiming to selectively erase specific generative capabilities while preserving the model’s ability to generate unrelated content—without requiring full retraining.
-Several existing methods have been developed for concept unlearning in T2I diffusion models.
-ESD [12] and CA [23] employ a frozen guiding model to provide negative guidance for fine-tuning.
-UCE [13] and RECE [14] leverages a closed-form solution to modify cross-attention weights, enabling simultaneous unlearning of multiple concepts.
-Salun [11] and Scissorhands [52] optimize parameter updates based on their influence on concept retention.
-Meanwhile, RACE [21] and AdvUnlearn [59]generate adversarial samples to enhance unlearning and improve model robustness.
-Although concept unlearning methods provide a certain level of security for T2I diffusion models, their effectiveness remains questionable when faced with diverse real-world scenarios and adversarial examples.
-Recent studies on adversarial attacks [51, 6, 58] against T2I DMs indicate that existing concept unlearning methods remain vulnerable to such attacks.
-To enhance robustness, current approaches rely on increasing the number of fine-tuning iterations or incorporating adversarial training [59, 21].
-However, these methods significantly increase computational cost while further degrading the model’s generative quality.
-Unlike previous approaches [23, 12, 53, 33], KSCU explicitly considers the relationship between the stepwise denoising process of diffusion models and the effectiveness of concept unlearning.
-Instead of modifying the model across all denoising steps, KSCU selectively applies unlearning to specific denoising steps that contribute most significantly to the target concept’s generation.
-This targeted approach reduces the number of training iterations required for unlearning, thereby minimizing unnecessary parameter adjustments and preserving the model’s overall generative quality.
-3 Key Step Concept Unlearning
-As shown in Figure 2, our framework consists of three main components: the Key Step Unlearning Loss module, including the unlearn model that undergoes fine-tuning and the guidance model that provides reverse guidance; the Key Step Table module, which controls the denoising steps during training; and the Prompt Augmentation module to help the model learn more generalized semantic distribution of the unlearned concept in the text embedding space.
-3.1 Background
-Diffusion models [17] are a class of generative models that learn data distributions by gradually denoising a normally distributed variable.
-The process consists of two phases: the forward diffusion process and the reverse denoising process.
-In the forward process, Gaussian noise is incrementally added to the data over a series of time steps, transforming the data into a nearly isotropic Gaussian distribution.
-Conversely, the reverse process learns to iteratively denoise the data, recovering the original distribution.
-Latent Diffusion Models.
-While traditional diffusion models operate directly in the high-dimensional data space, latent diffusion models (LDMs) [40] introduce a more efficient approach by performing the diffusion process in a compressed latent space.
-LDMs first employ an autoencoder to encode the input data $\mathbf{x}$ into a lower-dimensional latent representation $\mathbf{z}=E(\mathbf{x})$, where $E$ is the encoder.
-The diffusion process is then applied to $\mathbf{z}$ instead of $\mathbf{x}$, significantly reducing computational complexity.
-The key distinction of LDMs lies in their separation of the diffusion process from the high-dimensional data space. The forward process in the latent space is defined as:
-$q(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\mathbf{z}_{t};\sqrt{1-\beta_{t}}\mathbf{z}_{t-1},\beta_{t}\mathbf{I}),$ (1)
-where $\mathbf{z}_{t}$ is the latent representation at time step $t$. The reverse process learns to denoise the latent representation:
-$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},t),\Sigma_{\theta}(\mathbf{z}_{t},t)).$ (2)
-After the reverse process, the decoder $D$ maps the denoised latent representation $\mathbf{z}_{0}$ back to the data space: $\mathbf{x}=D(\mathbf{z}_{0})$.
-This two-stage approach not only reduces computational costs but also enables the generation of high-resolution samples by focusing on a compact latent space.
-Classifier-Free Guidance.
-Classifier-Free Guidance [16] is a technique designed to improve the controllability of diffusion models without relying on an external classifier.
-Instead of conditioning the model on explicit labels, classifier-free guidance jointly trains a conditional and an unconditional diffusion model.
-Formally, let $\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ denote the conditional prediction and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ denote the unconditional prediction, where $\mathbf{c}$ is the conditioning information (e.g., a class label or text embedding).
-The guided prediction is computed as:
-$\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))$ (3)
-where $w$ is the guidance scale controlling the strength of conditioning. A higher $w$ emphasizes the conditioning information, while $w=0$ reduces the model to an unconditional generator.
-3.2 Key Step Unlearning Loss
-Figure 3: Sampling with ”nudity woman” prompt reveals emergence of nudity concepts in the final 10 denoising steps.
-According to formula (3), in a classifier-free guidance diffusion model, the linear extrapolation of $\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ and $\epsilon_{\theta}(\mathbf{x}_{t},t)$ serves as the gradient of the implicit classifier $\nabla_{x_{t}}\log p(y|x_{t})$.
-Our objective is not to generate images corresponding to the text prompt $c$, but rather to shift the model’s predicted noise $\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ in the opposite direction of $\nabla_{x_{t}}\log p(y|x_{t})$.
-Furthermore, we exclusively fine-tune the conditional noise $\hat{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ to ensure that the generation of unrelated concepts remains unaffected.
-The fine-tuning objective is formulated as:
-$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)-w\cdot(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot(\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t))\end{split}$ (4)
-From this, we derive the unlearning loss:
-$\begin{split}\text{Loss}_{\text{unlearn}}&=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c})-(\epsilon_{\theta}(\mathbf{x}_{t},t)-\left(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{x}_{t},t)\right))\right\|_{2}\\ \end{split}$ (5)
-where $\epsilon^{*}_{\theta}$ represents the unlearned model, $\epsilon_{\theta}$ is the guidance model, $x_{t}$ denotes the denoised image at step $t$, and $c$ corresponds to the text prompt.
-More importantly, the Key Step Unlearning Loss also accounts for the optimization of unconditional noise.
-To preserve the model’s ability to generate unrelated concepts, we introduce a regularization term based on unconditional noise prediction to minimize parameter changes.
-Specifically, we constrain the L2 distance between the unconditional noise predictions of the guidance and unlearned models, offering partial protection for shared parameters.
-Additionally, we observe that unconditional noise, which uses the previous sampling step as a prior, gradually aligns with the target concept as denoising progresses.
-As shown in Figure 3, we use the prompt ”nudity woman” with DDIM steps set to 50 and sample images using Stable Diffusion v1-4.
-From left to right, the images correspond to step = 40, step = 50, and the last 10 steps sampled using only unconditional noise.
-The qualitative results show that even unconditional noise alone causes the generation to drift toward the ”nudity” concept.
-This highlights the necessity of applying unlearning not only to the conditional branch but also to the unconditional branch in classifier-free guidance diffusion models.
-Accordingly, we apply unlearning training to the unconditional component as well.
-The regularization loss is formulated as:
-$\text{Loss}_{\text{regular}}=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t)-((1+\lambda_{1})\epsilon_{\theta}(\mathbf{x}_{t},t)-\lambda_{1}\ \epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c}))\right\|^{2}_{2}$ (6)
-where $\lambda_{1}$ needs to be a small value that is linearly related to $t$. For $t$ from 1000 to 0, we set $\lambda_{1}=(1000-t)/(2*10^{5})$
-The final Key Step Unlearning Loss function for KSCU is defined as:
-$\text{Loss}=\text{Loss}_{\text{unlearn}}+\lambda_{2}\cdot\text{Loss}_{\text{regular}}$ (7)
-where $\lambda_{2}=1\times 10^{-4}$ in KSCU.
+When trained on such datasets, T2I diffusion models inevitably acquire the ability to reproduce these sensitive patterns, raising significant security and ethical concerns.
+To address this issue, concept unlearning has been introduced as a promising direction, aiming to selectively erase specific generative capabilities while preserving the model’s ability to synthesize unrelated content—without requiring costly full retraining.
+A variety of approaches have been proposed to realize this goal.
+Early methods such as ESD [29] and CA [22] rely on a frozen guiding model to provide negative supervision during fine-tuning, effectively discouraging the generation of undesired concepts while leaving unrelated knowledge largely intact.
+UCE [25] and RECE [26] instead adopt closed-form updates to cross-attention weights, enabling efficient unlearning of multiple concepts simultaneously.
+Parameter-significance-based approaches like Salun [23] and Scissorhands [48] update weights selectively according to their contribution to concept retention, minimizing interference with other learned capabilities.
+In parallel, adversarial strategies such as RACE [30] and AdvUnlearn [49] generate challenging counterexamples, which improve both the effectiveness and robustness of unlearning.
+More recently, SPEED [50] and MACE [51] have advanced scalable multi-concept unlearning: SPEED formulates erasure as a null-space constrained optimization for efficient removal while preserving unrelated priors, whereas MACE leverages cross-attention refinement and modular LoRA fusion to support large-scale erasure across hundreds of concepts.
+Despite these advances, existing methods provide only partial mitigation.
+Their reliability remains limited when confronted with diverse real-world scenarios or adversarial prompts.
+Recent studies on attacks against T2I diffusion models [52, 53, 54] further highlight that current unlearning techniques are vulnerable, underscoring the need for more effective and robust solutions.
+III Key Step Concept Unlearning
+As shown in Figure 2, our framework consists of three main components: the Key Step Table module, which controls the denoising steps during training; the Prompt Augmentation module to help the model learn a more generalized semantic distribution of the unlearned concept in the text embedding space; and the Key Step Unlearning Optimization module, including the unlearned model that undergoes fine-tuning and the guidance model that provides reverse guidance.
+III-A Motivation & Key Step Table Generation
+Figure 3: Correlation between text and attention map at different steps
+TABLE I: Preliminary Verification Experiment.
+The superior UA and UDA performance of ESD early 70% steps is attributed to model collapse.
+While ESD last 70% steps shows comparable unlearning effectiveness to full-step ESD, with less damage to the generative capability and improved time efficiency.
+[…]
+ESD 88.2 73.7 15.4 2557
+ESD last 70% steps 85.9 80.9 15.2 2072
+ESD early 70% steps 90.4 57.7 69.7 1355
+Due to the stepwise nature of diffusion-based generation, each sampling step contributes differently to the final image.
+Existing work indicates that earlier denoising steps predominantly capture low-frequency structures, while later steps progressively refine high-frequency details [33, 32].
+This leads to a layered emergence of semantic concepts during generation, as illustrated in Figure 3.
+Motivated by this observation, we hypothesize that concept unlearning can be achieved by modifying only the later denoising steps to adjust high-frequency information with greater semantic impact, replacing the target concept with a nearby but semantically distinct alternative (e.g., wolf $\rightarrow$ dog), thereby preventing the unwanted concept from appearing in the generated output.
+To validate this hypothesis, we conduct a preliminary experiment using the ESD method [29] on Stable Diffusion v1-4, restricting optimization to either the first or last 70% of denoising steps.
+As shown in Table I, training on the early 70% leads to model collapse, while fine-tuning the last 70%—with 30% fewer iterations—retains 95% of the full-step unlearning accuracy.
+These results indicate that targeting high-frequency denoising steps has the potential to achieve effective concept unlearning while better preserving the model’s generative capability with improved time efficiency.
+Based on the above conclusion, we design a step selection algorithm to generate Key Step Table.
+The Key Step Table consists of incrementally growing step intervals and shifts the starting point forward after a fixed number of loops to progressively cover a broader range of late denoising steps.
+As shown in Algorithm 1, the algorithm constructs the Key Step Table from a given starting step $S$ to an ending step $E$.
+In each iteration, the full interval from $s_{\text{cur}}$ to $E$ is appended to the Key Step Table.
+After $loop_{n}$ iterations, the starting step $s_{\text{cur}}$ is shifted forward by one, and the process is repeated until the Key Step Table reaches the desired length $L$.
+This mechanism ensures that training focuses more on later denoising steps, which are more influential, while avoiding overfitting to a single segment.
 Algorithm 1 Key Step Table Generation
 Input:start step $S$, end step $E$, table length $L$, full loops before shift $loop_{n}$
+Output:$KeyStepTable[:L]$
 […]
-1:  $KeyStepTable\leftarrow None$
-2:  $s_{cur}\leftarrow S$
-3:  $loop_{cur}\leftarrow 0$
+[…]
+[…]
 4:while $len(KeyStepTable)<L$ do
-5:     $KeyStepTable\leftarrow KeyStepTable\cup[s_{cur},s_{cur}+1,\dots,E]$
+5:  $KeyStepTable\leftarrow KeyStepTable\cup[s_{\text{cur}},s_{\text{cur}}+1,\dots,E]$
 6:  $loop_{cur}\leftarrow loop_{cur}+1$
 7:  if $loop_{cur}=loop_{n}$ then
-8:        $s_{cur}\leftarrow\min(s_{cur}+1,E-1)$
+8:   $s_{\text{cur}}\leftarrow\min(s_{\text{cur}}+1,E-1)$
 9:   $loop_{cur}\leftarrow 0$
@@
 11:end while
-3.3 Key Step Table Generation
-The goal of concept unlearning is to prevent the generation of a target concept.
-Due to the nature of diffusion-based generation, high-frequency details in the image are primarily refined during the later denoising steps.
-Therefore, we hypothesize that these later steps play a more critical role in the unlearning process.
-To validate this, we design a preliminary experiment using ESD [12] to fine-tune Stable Diffusion v1-4, while restricting the range of training steps.
-The results show that, for NSFW content unlearning, training only the last 70% of denoising steps (with the total number of training iterations also reduced to 70% of the original) preserves approximately 95% of the unlearning accuracy observed when training across all steps.
-This indicates that, compared to early steps, the later steps contribute more significantly to concept forgetting.
-Limiting unlearning to these steps—those responsible for high-frequency detail generation—can still achieve effective unlearning performance.
-Based on the above observation, we design a step selection algorithm to generate Key Step Table.
-The Key Step Table consists of incrementally growing step intervals and shifts the starting point forward after a fixed number of loops to progressively cover a broader range of late denoising steps.
-As shown in Algorithm 1, the algorithm constructs the Key Step Table from a given starting step $S$ to an ending step $E$.
-In each iteration, the full interval from $s_{cur}$ to $E$ is appended to the Key Step Table.
-After $loop_{n}$ iterations, the starting step $s_{cur}$ is shifted forward by one, and the process is repeated until the Key Step Table reaches the desired length $L$.
-This mechanism ensures that training focuses more on later denoising steps, which are more influential, while avoiding overfitting to a single segment.
-Furthermore, our analysis indicates that the most influential steps vary by concept type.
-Specifically, for NSFW content and class unlearning, the last 70% of denoising steps are most critical, whereas for style unlearning, only the last 50% of steps play a dominant role, while for individual instance forgetting, this interval can be further reduced to the last 20%.
-These observations highlight the need for a structured step selection approach.
-3.4 Prompt Augmentation
-Figure 4: Comparison of Semantic Range in Prompt Embedding Space.
-Left: Limited effective unlearning semantic range (red dashed boundary) with single target concept prompts.
-Right: Expanded effective unlearning semantic range (green dashed boundary) using augmented prompts.
-For prompt-based concept unlearning in T2I DMs, the objective is to erase undesired visual concepts from the model’s generated results.
-Prompts describe these visual concepts to explicitly specify the target concept for unlearning.
-However, accurately depicting a complex visual concept through textual prompts is often challenging.
-Inaccurate prompts may lead to incomplete unlearning, where related concepts remain, or incorrect removal of unrelated concepts.
-Recently, many methods [54, 5] have recognized this issue and introduced improvements targeting prompt section.
-In this section, we propose a simpler yet effective Prompt Augmentation method to address this problem.
-For a given concept $c$ to be unlearned, we use large language models such as ChatGPT [32] and DeepSeek [8] to generate ten augmentation rules that expand the original concept into diverse textual descriptions.
-For example, given the concept ”cat” and the rule ”a picture of {}, we obtain: ”a picture of a cat”.
-We then introduce perturbations by randomly inserting scrambled characters, shuffling word order, or removing non-keywords.
-Additionally, Gaussian noise is injected into the text embeddings of these modified descriptions [19].
-Since the embeddings are ultimately perturbed, the quality of the augmentation rules need not be strictly controlled.
-As shown in Figure 4, in the embedding space [36], nearby vectors share similar semantics and induce similar generative behavior.
-Leveraging this property, Prompt Augmentation enables the model to better perceive the semantic space of the target concept during unlearning, leading to more effective concept erasure.
-Theoretically, Prompt Augmentation constructs numerous embeddings close to the target concept, aligning with its semantic distribution.
-This allows the model to unlearn not only the concept itself but also its semantic variations.
-In contrast, adversarial samples are discrete and harder to construct.
-Compared to adversarial training, which is costly in time and computation, Prompt Augmentation offers a more efficient alternative for enhancing unlearning robustness.
-4 Experimental Results
-Table 1: Class and Style Unlearning Performance on Unlearn Canvas. The bold font represents the best, and the underlined font represents the second-best.
-Method class style
+III-B Prompt Augmentation
+In prompt-based concept unlearning for T2I diffusion models, the target visual concept is specified via text prompts to guide the model in identifying what to forget.
+However, accurately capturing the semantics of a complex visual concept through textual descriptions is often non-trivial.
+For instance, some concepts may have multiple contextual meanings, and vague or underspecified prompts may lead to incomplete unlearning, where related sub-concepts persist, or to overgeneralized forgetting, where unrelated concepts are erroneously removed.
+Recent methods [55, 56] have recognized this limitation and introduced enhancements to the prompt design stage, often by refining textual formulations or aligning prompts with concept embeddings.
+To address this problem from a more lightweight perspective, we propose a simple yet effective Prompt Augmentation module.
+For a given concept $c$ to be removed, our method first leverages large language models (e.g., ChatGPT [57] and DeepSeek [58]) to automatically generate a set of augmentation rules, expanding the original prompt into a diverse collection of textual variants (e.g., “An augmented prompt about the {target concept}).
+In practice, we typically sample ten such rules, which cover lexical, syntactic, and contextual variations of the target concept.
+To further enhance robustness, we incorporate additional augmentation strategies: injecting random character noise at the beginning or end of prompts (to simulate adversarial prefix/suffix attacks commonly observed in safety evaluations), shuffling word order to test semantic invariance, and removing non-essential words to reduce overfitting to specific phrasings.
+Beyond the textual space, we also perturb the prompt embeddings directly by adding Gaussian noise, thereby generating a dense set of representations around the target concept in the embedding space.
+This forces the model to learn to suppress the target concept across a continuous semantic neighborhood rather than a few discrete prompt formulations.
+Importantly, since the embeddings are ultimately perturbed, the strict quality of the generated textual rules is less critical—what matters is the diversity and coverage of the augmented distribution.
+Overall, Prompt Augmentation provides a lightweight yet powerful mechanism to mitigate the brittleness of text-prompt supervision in concept unlearning.
+By diversifying both textual and embedding-level representations, it equips the unlearning process with stronger resistance to adversarial prompts and improves generalization to real-world usage scenarios.
+III-C Key Step Unlearning Optimization
+Figure 4: Image generation with Stable Diffusion v1-4 using the prompt “naked woman”: evolution of nude scores and qualitative results across the denoising process.
+Conditional Predicted Noise Optimization.
+According to formula (3), in a CFG diffusion model, the linear extrapolation of $\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ and $\epsilon_{\theta}(\mathbf{z}_{t},t)$ serves as the gradient of the implicit classifier $\nabla_{z_{t}}\log p(y|z_{t})$.
+Our objective is not to generate images corresponding to the text prompt $c$, but rather to shift the model’s predicted noise $\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ in the opposite direction of $\nabla_{z_{t}}\log p(y|z_{t})$.
+Furthermore, we exclusively fine-tune the conditional noise $\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ to ensure that the generation of unrelated concepts remains unaffected.
+The fine-tuning objective is formulated as:
+$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{z}_{t},t)-w\cdot(\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{z}_{t},t))\\ &=\epsilon_{\theta}(\mathbf{z}_{t},t)+w\cdot(\epsilon_{\theta}^{*}(\mathbf{z}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{z}_{t},t))\end{split}$ (4)
+From this, we derive the unlearning loss like ESD [29]:
 […]
-ESD [12] 68.6 98.9 96.4 26.1 100.0 85.7 99.0 25.3
-CA [23] 32.5 86.7 83.4 51.9 96.0 86.6 88.7 37.4
-FMN [56] 25.9 93.0 96.1 36.8 41.5 95.1 91.3 28.5
-EDiff [53] 78.0 94.1 90.1 48.2 67.5 93.7 98.3 39.8
-Salun [11] 51.3 93.5 94.6 47.9 17.5 96.3 95.2 34.8
-SHS [52] 62.5 54.6 56.7 79.4 65.0 51.1 29.4 65.9
-KSCU 70.0 99.0 97.1 25.8 100.0 88.7 99.3 23.1
-4.1 Experimental Setup
-Dataset.
-To rigorously evaluate various concept unlearning methods, we conducted experiments on four concept unlearning task types: Class, Style, NSFW, and Instance on two datasets, including the I2P [44] and Unlearn Canvas [57].
-I2P is a dataset containing 4,703 NSFW prompts.
-For our experiments, we selected 931 prompts labeled as ”sexual” to generate images.
-We then used Nudenet [34] (threshold = 0) to analyze the generated images, detecting 575 nudity regions in the outputs from these 931 prompts.
-Across the entire I2P dataset, Nudenet identified a total of 793 nudity regions.
-This dataset setting reduced the computational cost while still preserving the majority of the experimental requirements.
-Unlearn Canvas is a recently released dataset comprising 20 classes and 60 distinct styles of high-resolution stylized images.
-Since testing the entire dataset was computationally intensive, we selected 10 classes and 10 styles for the evaluation, including the following:
-  • Class: Architectures, Bears, Cats, Dogs, Humans, Towers, Flowers, Sandwiches, Trees, Sea.
+where $\epsilon^{*}_{\theta}$ represents the unlearned model, $\epsilon_{\theta}$ is the guidance model, $z_{t}$ denotes the denoised data at step $t$, and $c$ corresponds to the text prompt.
+Unconditional Predicted Noise Optimization.
+More importantly, to the best of our knowledge, this is the first work to incorporate unconditional noise prediction into the optimization scope of concept unlearning.
+Initially, our goal was to constrain changes in the model’s unconditional prediction during training, aiming to preserve its generative performance.
+However, we identified a critical issue: although the prior of unconditional noise prediction is only conditioned on $\mathbf{z}_{t}$ and $t$, when $\mathbf{z}_{t}$ contains sufficient semantic information about the target concept $c$, the model’s unconditional prediction inevitably drifts toward $c$.
+To verify this, we empirically perform sampling using the prompt “naked woman”.
+At different denoising steps, we switch to purely unconditional noise prediction for the remaining steps, and apply NudeNet [59] to detect and score nudity in each generated image.
+As shown in Figure 4, even switching to unconditional sampling after just 5 standard steps produces images containing explicit nudity content.
+Furthermore, when switching after 30 steps, the generated results are almost indistinguishable from those produced by the full 50-step sampling.
+This indicates that, as the reverse process progresses, the sampling direction of unconditional noise prediction gradually shifts toward the concept $c$.
+Therefore, we argue that unconditional prediction should also be partially optimized for concept unlearning. We propose the following regularization loss:
+$\text{Loss}_{\text{regular}}=\left\|\epsilon_{\theta}^{*}(\mathbf{z}_{t},t)-\left((1+\lambda_{1})\epsilon_{\theta}(\mathbf{z}_{t},t)-\lambda_{1}\,\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\right)\right\|_{2}^{2}$ (6)
+Here, $\lambda_{1}$ is designed as a step-dependent coefficient to regulate the extent of concept unlearning applied to unconditional noise prediction.
+Since the bias toward concept $c$ increases with the number of reverse steps, we define $\lambda_{1}$ as a linear function with respect to $t$.
+Specifically, for $t$ ranging from 1000 to 0, we set $\lambda_{1}=(1000-t)*(2\times 10^{-5})$.
+The value was chosen empirically to balance semantic retention in early steps and suppression in later ones.
+The final Key Step Unlearning Optimization function for KSCU is defined as:
+$\text{Loss}=\text{Loss}_{\text{unlearn}}+\lambda_{2}\cdot\text{Loss}_{\text{regular}}$ (7)
+where $\lambda_{2}=1\times 10^{-4}$ in KSCU.
+III-D Training Algorithms
+Algorithm 2 KSCU-Quick
+Input: Unlearn Model $\epsilon^{*}_{\theta}$, Guidance Model $\epsilon_{\theta}$, scheduler $S$, target concept $c$, augmentation steps $N$, training steps $M$, Key Step Table
+Output: $\epsilon^{*}_{\theta}$
+1: Sample noise $n\sim\mathcal{N}(0,1)$
+[…]
+[…]
+4:for $i=2$ to $M$ do
+5:  if $i>N$ then
+6:   $c^{*}\leftarrow\text{augment}(c)$
+7:  else
+8:   $c^{*}\leftarrow c$
+9:  end if
+[…]
+11:  if $i<N$ then
+12:   $t_{next}$ = $KeyStepTable[i+1]$
+13:   if $t_{next}>t$ then
+14:    $z_{t_{next}}=D(\epsilon^{*}_{\theta},z_{t},t_{next},t,c)$
+15:   else
+16:    $z_{t_{next}}=D(\epsilon^{*}_{\theta},n,t_{next},0,c)$
+17:   end if
+18:   $z_{t}=z_{t_{next}}$
+19:  end if
+20:end for
+Although our KSCU framework significantly reduces the number of training iterations compared with conventional concept erasure approaches, the improvement in wall-clock training time remains limited.
+This is mainly because KSCU requires generating unlearning samples on-the-fly during fine-tuning, rather than pre-constructing the dataset as in other methods.
+Nevertheless, we argue that this online data generation strategy is crucial for the effectiveness of KSCU and thus remains an integral part of our design.
+On the other hand, we further observed that, benefiting from the step selection strategy in Algorithm 1, the steps chosen in the KeyStepTable exhibit a periodic and incrementally increasing pattern.
+Inspired by this observation, we propose an acceleration strategy that leverages such regularity to skip redundant intermediate sampling steps and only perform noise propagation and state updates at critical positions.
+This design substantially speeds up the training process while maintaining comparable performance to the original method.
+Concretely, the acceleration strategy reuses forward trajectories and recomputes states only at periodic checkpoints, effectively eliminating the computational overhead of repeated step-wise sampling.
+The complete KSCU algorithm, including this acceleration strategy, is summarized in Algorithm 2.
+IV Experimental Results
+TABLE II: Class and Style Unlearning Quantitative Performance on Unlearn Canvas.
+class style
+[…]
+[…]
+ESD [29] 0.81 68.6 98.9 96.4 26.1 0.91 100.0 85.7 99.0 25.3
+RECE [26] 0.70 62.8 88.9 96.3 25.9 0.00 12.0 97.6 98.4 20.9
+DUO [60] 0.00 27.8 89.9 75.2 95.8 0.00 31.5 77.0 89.8 69.1
+CA [22] 0.08 32.5 86.7 83.4 51.9 0.63 96.0 86.6 88.7 37.4
+FMN [24] 0.00 25.9 93.0 96.1 36.8 0.28 41.5 95.1 91.3 28.5
+ANT [61] 0.21 44.1 80.7 89.6 54.7 0.01 14.5 91.7 78.2 45.6
+EDiff [27] 0.68 78.0 94.1 90.1 48.2 0.38 67.5 93.7 98.3 39.8
+Salun [23] 0.33 51.3 93.5 94.6 47.9 0.04 17.5 96.3 95.2 34.8
+SHS [48] 0.16 62.5 54.6 56.7 79.4 0.04 65.0 51.1 29.4 65.9
+KSCU 0.84 70.0 99.0 97.1 25.8 0.95 100.0 88.7 99.3 23.1
+IV-A Experimental Setup
+Dataset. We evaluate concept unlearning across four task types—Class, Style, Instance, and Nudity—on two datasets: I2P [62] and Unlearn Canvas [63].
+I2P contains 4,703 NSFW prompts.
+We selected 931 prompts labeled as “sexual” to generate images, from which Nudenet [59] (threshold = 0.2) detected 575 nudity regions.
+Across the full dataset, 793 nudity regions were found.
+This subset retains most of the evaluation potential while reducing computational cost.
+Unlearn Canvas is a high-resolution stylized image dataset with 20 classes and 60 styles. Due to the high cost of full evaluation, we selected 10 diverse classes and 10 representative styles:
+  • Classes: Architectures, Bears, Cats, Dogs, Humans, Towers, Flowers, Sandwiches, Trees, Sea.
   • Styles: Abstractionism, Artist Sketch, Cartoon, Impressionism, Monet, Pastel, Pencil Drawing, Picasso, Sketch, Van Gogh.
-Such the selection also ensured diversity: the chosen classes include animals, plants, natural scenes, and man-made objects, while the styles cover both artistic styles such as ”Monet” and illustrative styles such as ”Sketch”.
+Such selections also ensured diversity: the chosen classes include animals, plants, natural scenes, and man-made objects, while the styles cover both artistic styles, such as “Monet,” and illustrative styles, such as “Sketch”.
 Evaluation.
-In our experiments, we evaluated the unlearning effectiveness of various methods using Unlearn Accuracy (UA) originally from [12].
-To achieve a more concise and effective evaluation of unlearning effectiveness, we computed UA as follows: $UA=1-x/y$.
-In the I2P dataset, $x$ represented the number of nudity regions detected in the output images by the unlearned model, and $y$ was 575.
-Compared to previous methods [12] that evaluated each exposed region individually, we aggregated all cases into a single metric, Unlearn Accuracy, which offered a more concise, analyzable and convincing measure of unlearning performance.
-For the Unlearn Canvas dataset, $x$ denoted the number of target concepts detected by the classifier [10], while $y$ corresponded to the total number of generated images based on the target prompts.
-We also utilized the Fréchet Inception Distance (FID) [15], a widely used metric for assessing the quality of generated images, which measured the distance between the generated and real images in feature space.
-In the I2P dataset, we generated images from corresponding prompts using the COCO [24] dataset and computed the FID by comparing them with real images from coco30k to assess how well each baseline preserved the model’s generative capacity.
-For Unlearn Canvas, we generated images using the prompt ”A {class} image in {theme} style.”, then calculated the FID by comparing these images with real images from the Unlearn Canvas dataset.
-Besides UA and FID, Unlearn Canvas introduced two additional metrics: In-domain Retain Accuracy (IRA) and Cross-domain Retain Accuracy (CRA), which helped evaluate the model’s ability to retain relevant information.
-Lastly, we assessed the model’s defense capabilities by constructing adversarial text prompts using the P4D [6] and Unlearndiff(UDA) [58] methods based on I2P prompts, comparing them using Attack Success Rate (ASR).
-Figure 5: The figure demonstrates KSCU’s unlearning effectiveness and generative retainability on Unlearn Canvas.
-Images within the red border show the removal of target concepts, while images outside the red border highlight KSCU’s generative retainability.
-KSCU only aims to remove the target concept without enforcing constraints on the final output.
-In Appendix, we provide a controllable version of KSCU that can replace the target concept with any specified alternative concept.
-Baseline and Training Details.
-To provide a comprehensive comparison of various methods, we selected six state-of-the-art (SOTA) unlearning techniques, including ESD [12], CA [23], FMN [56], EDiff [53], Salun [11], and SHS [52].
-For KSCU’s training setup, we used a batch size of 1.
-The scheduler was set to 50 timesteps.
-For style unlearning tasks, the Key Step Table was configured to start from step 25 and end at step 50.
-For other tasks, the training Key Step Table began at step 15.
-We employed the Adam optimizer with a learning rate of $1\times 10^{-5}$, applying a total of 750 parameter updates.
-For all baseline methods, we adopted the default hyperparameters as specified in their original papers.
-We employed DDIM for denoising, using 50 steps.
-For Unlearn Canvas, we used Stable Diffusion 1.5, while for all other experiments, unless explicitly stated otherwise, Stable Diffusion 1.4 was used.
+We adopt Unlearn Accuracy (UA) [29] as the evaluation metric, computed as $\text{UA}=1-x/y$.
+In I2P, $x$ is the number of nudity regions detected post-unlearning, and $y=575$.
+Unlike prior methods [29] that report the number of nude parts, we aggregate all detections into a single UA score for clarity and comparability.
+In Unlearn Canvas, $x$ is the number of classifier-detected target concepts, and $y$ is the number of generated images for the corresponding prompts.
+We also report Fréchet Inception Distance (FID) [64] to measure image quality.
+For I2P, prompts are drawn from COCO [65], and FID is computed against real images from coco30k.
+For Unlearn Canvas, images are generated with the prompt “A {class} image in {theme} style,” and compared to real images in the dataset.
+For Unlearn Canvas, we further report In-domain Retain Accuracy (IRA) and Cross-domain Retain Accuracy (CRA) to assess the preservation of non-target information.
+To evaluate defense performance, we generate adversarial prompts using UDA [54], reporting Attack Success Rate (ASR) on I2P.
+More importantly, to intuitively compare the overall performance of different methods in terms of unlearning effectiveness and generative retainability, we introduce a statistical metric, $\text{Overall}=\text{UA}_{0\text{-}1}\times(1-\text{FID}_{0\text{-}1}),$ where $\text{UA}_{0\text{-}1}$ and $\text{FID}_{0\text{-}1}$ denote the min-max normalized UA and FID.
+Only methods that achieve both effective unlearning and high-generative retainability yield high overall scores, making it a direct indicator of performance trade-off.
+Figure 5: KSCU’s Qualitative Performance on Unlearn Canvas.
+Red-bordered images show the generations of the unlearned concept after KSCU, while borderless images depict the results of unrelated concepts.
+More qualitative results and controlled replacement experiments are provided in Appendix A.
+Baselines and Training Details.
+We compare KSCU with nine SOTA methods: ESD [29], RECE [26], CA [22], FMN [24], ANT [61] DUO [60], EDiff [27], Salun [23], and SHS [48].
+All baselines use default hyperparameters from their original papers.
+KSCU was trained with a batch size of 1 and 50 timesteps.
+For style unlearning, the Key Step Table started at $S=25$ with $loop_{n}=8$, updating only the cross-attention modules for 500 optimizer iterations.
+For other tasks, it started at $S=15$ with $loop_{n}=8$, updating all modules except cross-attention in the U-Net for 700 iterations.
+We used the Adam optimizer with a learning rate of $1\times 10^{-5}$.
+All experiments employed DDIM with 50 denoising steps.
+Stable Diffusion v1.5 was used for Unlearn Canvas, while all other experiments (unless otherwise specified) used v1.4.
+Results for more versions of SD are included in the Appendix.
 All experiments were conducted on a single NVIDIA L40 GPU.
-4.2 Class and Style Unlearning
+IV-B Class and Style Unlearning
 To evaluate the effectiveness of our proposed KSCU, we first performed qualitative experiments on class and style unlearning tasks.
-As shown in Figure 5, for different image samples, KSCU has completely blocked the generation of the target class or style, with minimal impact on unrelated classes and styles.
-Then we quantitatively evaluated its performance, as described below.
+As shown in Figure 5, for different image samples, KSCU has completely blocked the generation of the target class or style (red-bordered images), with minimal impact on unrelated classes and styles (borderless ones).
+We then quantitatively evaluate its performance as follows.
 Class Unlearning.
-In the selection of test categories for class unlearning, we randomly selected ten classes encompassing a diverse range from Unlearn Canvas, including animals, plants, man-made objects, and natural elements.
-The experimental results presented in Table 1 indicated that KSCU achieved superior category-wide unlearning while better preserving the model’s generative capability—an advantage not observed in other SOTAs.
-Specifically, as illustrated in Table 1, KSCU achieved 70%, 99.0%, 97.1%, and 25.8 on UA, IRA, CRA, and FID, respectively.
+As shown in Table II, KSCU achieves the highest overall score, indicating superior class unlearning while better preserving generative capability—an advantage not observed in other SOTA methods.
+Specifically, as illustrated in Table II, KSCU achieved 70%, 99.0%, 97.1%, and 25.8 on UA, IRA, CRA, and FID, respectively.
 In terms of UA, KSCU ranked as the second-best method, while it achieved the best performance on IRA, CRA, and FID.
@@
 Conversely, while EDiff attained the highest UA, its FID reached 48, suggesting that although it exhibited strong unlearning performance, it significantly degraded the model’s generative ability.
-This low generative retainability of EDiff indicates that it is difficult to use in practical applications.
+The low generative retainability of EDiff makes it difficult to use in practical applications.
 Style Unlearning.
-We conducted experiments on image style unlearning using the Unlearn Canvas dataset.
-We quantitatively assessed the feasibility of removing artistic styles to address copyright concerns using the IRA and CRA metrics.
-Given computational constraints, we selected 10 styles from Section 4.1, covering three artistic styles and seven artistic representations.
-This selection ensured a diverse evaluation of style unlearning across different visual attributes while maintaining a feasible training cost.
-The experimental results presented in Table 1 indicated that KSCU achieved strong unlearning effectiveness and retainability, while maintaining a favorable balance between the two on the style unlearning task.
-Specifically, as shown in Table 1, KSCU achieved scores of 100%, 88.7%, 99.3%, and 23.1 in UA, IRA, CRA, and FID, respectively.
-Among these, UA, CRA, and FID were the best across all compared methods.
-While KSCU’s IRA was slightly lower than the Salun method with very low unlearning effectiveness (UA $=$ 17.5%), this method reached high IRA at the expense of insufficient concept forgetting.
-In contrast, when compared with other high unlearning effectiveness methods such as CA and ESD (UA $>$ 95%), KSCU demonstrated clearly superior IRA, CRA, and FID.
-These results indicated that KSCU not only effectively eliminated traces of artistic styles but also better preserved the model’s ability to generate unrelated concepts.
-KSCU halved the iteration count compared to ESD by selectively fine-tuning only the latter 50% of denoising steps.
-Despite reducing the number of parameter updates by half,, KSCU still outperformed ESD in style unlearning, reinforcing that full-step fine-tuning is unnecessary for concept unlearning in T2I diffusion models.
-Focusing on the most influential steps enables more efficient and controlled unlearning while minimizing disruptions to the model’s generative capacity.
-These findings suggest that the Key Step Table optimization strategy adopted by KSCU offers a more practical and scalable solution for concept unlearning in large-scale diffusion models.
-Table 2: Evaluation of Unlearning Capability on ”Nudity”. The bold font represents the best, and the underlined font represents the second-best.
-Method UA(%$\uparrow$ P4D(%$\downarrow$ UDA(%$\downarrow$ FID-30$\downarrow$
-ESD [12] 88.2 39.3 73.7 15.4
-CA [23] 34.3 - - 15.3
-Salun [11] 100.0 0.0 4.2 67.9
-FMN [56] 38.6 - - 20.9
-EDiff [53] 93.9 29.6 52.5 19.8
-SHS [52] 84.3 62.2 95.5 18.1
-KSCU 96.5 23.2 47.2 14.1
-4.3 NSFW Unlearning
-NSFW Unlearning.
-In the NSFW concept forgetting experiment, we fine-tuned the models using ”nudity” as the target concept and obtained the corresponding unlearned models.
-Each method used its corresponding unlearned model to generate 931 and 30,000 images based on prompts from the i2p [44] and COCO [24] datasets, respectively.
-Table 2 presents the results, assessing the unlearning effectiveness of various methods using Unlearn Accuracy, P4D, and UDA’s ASR.
-FID is also compared to evaluate generative retainability.
-Note that CA and FMN, which fine-tune only the cross-attention module, show limited NSFW unlearning effects and are not further tested for robustness against attacks.
-As shown in Table 2, KSCU balanced unlearning effectiveness and generative retainability, achieving a UA of 96.5%, surpassing other methods in forgetting capability.
-Meanwhile, its FID-30K of 14.1 was the lowest among all baselines, confirming that KSCU best generative retainability.
-In contrast, although Salun achieves the highest UA, its FID of 67.9 reveals a significant degradation in generative quality.
-Moreover, we observed extensive large-area distortions in Salun’s generated results, indicating severe over-unlearning.
-Compared to methods with similar FID-30K scores, such as ESD and CA, KSCU demonstrated superior forgetting performance.
+As shown in Table II, KSCU demonstrates strong unlearning effectiveness and retainability on the style unlearning task, achieving the highest overall score of 0.95.
+Specifically, KSCU achieves scores of 100%, 88.7%, 99.3%, and 23.1 in UA, IRA, CRA, and FID respectively, as shown in Table II.
+Notably, KSCU achieves the best performance in UA and CRA among all compared methods.
+While its FID and IRA are slightly lower than those of RECE, the latter exhibits extremely poor unlearning efficiency (UA == 12.0%), achieving higher IRA or FID at the cost of insufficient concept forgetting.
+In contrast, compared to high-efficiency unlearning methods like ESD (UA >> 95%), KSCU achieves significantly better results in IRA, CRA, and FID.
+These results suggest that KSCU not only effectively removes stylistic traces but also better preserves the model’s ability to generate unrelated concepts.
+By selectively fine-tuning only the last 50% of the denoising steps, KSCU reduces the training iterations by half compared to ESD.
+Despite the substantial reduction in training cost, KSCU still outperforms ESD in style unlearning, further demonstrating that full-step fine-tuning is unnecessary for concept unlearning in T2I diffusion models.
+Focusing on the most influential steps enables more efficient and controllable unlearning while minimizing disruption to the model’s generative capabilities.
+These findings indicate that the key step table optimization strategy adopted by KSCU offers a more practical and scalable solution for concept unlearning in large-scale diffusion models.
+TABLE III: Evaluation of Unlearning Capability on “Nudity”.
+The bold font represents the best, and the underlined font represents the second-best.
+Time is the first run time.
+Some methods include time to generate the required data.
+[…]
+[…]
+ESD 0.71 88.2 73.7 15.4 2557
+RECE 0.66 85.4 87.5 15.0 22
+Salun 0.00 100.0 4.2 67.9 581
+DUO 0.00 56.2 98.3 15.2 628
+ANT 0.78 99.9 5.3 25.4 3016
+EDiff 0.77 93.9 52.5 19.8 1914
+SHS 0.59 84.3 95.5 18.1 817
+KSCU 0.92 96.5 47.2 14.1 546
+IV-C Nudity Unlearning & Robustness
+Figure 6: Comparison of visualization results of forgetting four groups of instances using various concept unlearning methods.
+It can be observed that our method(KSCU) not only achieves more thorough instance forgetting but also better preserves unrelated generative quality.
+Nudity Unlearning.
+In the NSFW concept unlearning experiment, we fine-tuned models to forget the concept of “nudity” and generated unlearned models.
+Each method was evaluated by generating 931 and 30,000 images using prompts from the i2p [62] and COCO [65] datasets, respectively.
+We compare unlearning effectiveness using Unlearn Accuracy and ASR, and assess generative retainability with FID.
+We also report the training time of each method to evaluate efficiency.
+As shown in Table III, KSCU achieves a overall score of 0.92 in the nudity unlearning task, significantly outperforming all other methods.
+This indicates a strong balance between unlearning effectiveness and generative retainability.
+Specifically, KSCU attains a UA of 96.5%, surpassing most baselines in forgetting capability.
+At the same time, its FID-30K reaches 14.1—the lowest among all compared methods—demonstrating KSCU’s superior generative retainability.
+In contrast, although Salun and ANT achieve the highest UA, its FID of 67.9 reveals a significant degradation in generative quality.
+Moreover, we observed extensive large-area distortions in their generated results, indicating severe over-unlearning.
+Compared to methods with similar FID-30K scores, such as ESD and RECE, KSCU demonstrated superior forgetting performance.
 Conversely, while methods like EDiff and Salun matched KSCU in unlearning effectiveness, they fell short in retaining generative ability.
-These results established KSCU as the most effective and balanced approach for NSFW concept unlearning.
+These results established KSCU as the most effective and balanced approach for nudity unlearning, and also the fastest among all fine-tuning-based methods in terms of training efficiency.
 Robustness.
-In the adversarial sample defense experiment, excluding Salun (which exhibited over-unlearning), KSCU achieved the lowest ASR against P4d [6] and UDA [58] among all other SOTAs, demonstrating its strong robustness.
-On the other hand, we observed a certain degree of correlation between the performance of different methods under adversarial attacks and their forgetting precision.
-We believed this was because the i2p prompts used in the experiment were manually selected, diverse prompts with inherent adversarial properties.
+In the adversarial sample defense experiment, excluding Salun and ANT (which exhibited over-unlearning), KSCU achieved the lowest ASR against UDA [54] among all other SOTAs, demonstrating its strong robustness.
+We also observed a certain degree of correlation between the performance of different methods under adversarial attacks and their forgetting precision.
+We believed this was because the i2p prompts used in the experiment were selected, diverse prompts with inherent adversarial properties.
 As a result, the unlearning performance on i2p can, to some extent, reflect the robustness of a given method in complex real-world scenarios or against adversarial samples.
-4.4 Instance Unlearning
-Figure 6: KSCU vs.
-ESD in Instance Unlearning.
-KSCU achieves forgetting with only 20% of ESD’s iterations.
-Despite ESD’s greater adjustments, KSCU demonstrates superior performance in erasing instance information.
-To demonstrate KSCU’s effectiveness in unlearning instances memorized by diffusion models, we qualitatively compared the unlearning performance of KSCU and ESD through visualized instance unlearning results.
-The reason for choosing ESD was that the architecture of ESD was more similar to ours and ESD had shown good performance in previous experiments.
-As shown in Figure 6, KSCU demonstrated superior performance in unlearning public figures.
-In KSCU’s results, facial features of public figures were effectively altered, whereas ESD partially retained the original characteristics.
-For artificial objects, ESD’s unlearning effect was barely noticeable, whereas KSCU replaced the target with either a visually similar but semantically distinct object (e.g., Eiffel Tower $\rightarrow$ highway) or a semantically related alternative (e.g., Titanic $\rightarrow$ small boat).
-KSCU, due to its well-designed approach, achieved a more effective instance unlearning performance.
-More importantly, KSCU achieved this with lower computational cost, highlighting its potential for broader applications.
-4.5 Ablation Study
-Table 3: Ablation Study: Impact of Prompt Augmentation(PA) and Key Step Table (with Different Start $S$) on KSCU’s Performance
-Method UA(%$\uparrow$ P4D(%$\downarrow$ UDA(%$\downarrow$ FID-30$\downarrow$
-ESD [12] 88.2 39.3 73.7 15.4
-KSCU w/o PA 92.5 30.3 51.8 14.5
-KSCU $S$=20 79.7 59.2 85.3 15.7
-KSCU $S$=10 96.1 23.0 48.1 14.3
-KSCU $S$=0 96.0 26.7 54.5 18.8
-KSCU 96.5 23.2 47.2 14.1
-In the ablation study, we mainly evaluated the impact of the Key Step Table and Prompt Augmentation.
-After removing Prompt Augmentation, KSCU exhibited a decrease in UA and a significant increase in ASR against adversarial examples, indicating a reduction in model robustness.
-This strongly confirmed that Prompt Augmentation enhanced the model’s ability to better perceive the semantic space of the target concept, achieve precise concept erasure, and improve robustness.
-In the ablation experiment of the Key Step Table’s start, we observed that when $S$=10 (covering 80% of steps), the performance was similar to $S$=15 (covering 70% of steps).
-However, when $S$=20 (covering 60% of steps), the unlearning effect of KSCU significantly decreased, suggesting that steps beyond 15 played a key role in the imaging of “nudity.” In the $S$=0 (covering 100% of steps) experiment, the unlearning performance remained strong.
-However, since all steps were affected, the FID increased, indicating a more significant negative impact on the model’s generative ability.
-Notably, the unlearning effectiveness of $S$=0 (100% steps), $S$=10 (80% steps), and $S$=15 (KSCU) was highly similar.
-This suggested that the first 30% of steps had little influence on the final rendering of ”nudity,” making their unlearning unnecessary.
-The results of the ablation experiment further demonstrated that for concept unlearning, KSCU, which focuses on unlearning key steps, was superior to the previous method of adjusting parameters for all steps.
-5 Conclusion
-This study investigates the impact of denoising steps on concept unlearning in diffusion models and proposes KSCU, a method that targets key denoising steps for concept unlearning.
-Through experiments on four different concept unlearning tasks, we demonstrate that KSCU surpasses previous state-of-the-art methods in terms of unlearning effectiveness, robustness, and generative retainability.
-The core ideas behind the key steps introduced in this work provide valuable insights for research in related fields.
-Appendix
+IV-D Instance Unlearning
+To further demonstrate the effectiveness of KSCU in unlearning specific instances memorized by diffusion models, we qualitatively compared KSCU with several state-of-the-art approaches using visualized instance unlearning results.
+The comparison provides a more holistic view of the trade-off between forgetting accuracy and generative retainability.
+The experiments reveal that, compared to other methods, KSCU not only removes the target instance more effectively but also achieves superior preservation of unrelated generative capabilities, thereby striking a better balance between forgetting and retention.
+As illustrated in Figure 6, KSCU consistently excels at unlearning public figures.
+In KSCU’s outputs, the background, lighting, and body posture remain largely unchanged, while the facial features of the public figure are effectively altered or replaced, thus achieving successful instance unlearning without degrading overall scene integrity.
+By contrast, RECE is able to suppress the identity of the public figure but does so at the cost of severely disrupting unrelated visual content, leading to unnatural or distorted backgrounds.
+DUO and ANT provide only limited suppression of the target instance, often leaving recognizable traces of the original identity, while simultaneously introducing distributional shifts that reduce the fidelity and stylistic consistency of the generated image.
+ESD, on the other hand, shows relatively strong preservation of unrelated content but fails to sufficiently erase instance-level information, resulting in outputs that still closely resemble the original subject.
+The comparison is even more striking in the case of landmark unlearning.
+DUO and ESD exhibit negligible forgetting effects, essentially leaving the landmark unchanged.
+RECE aggressively removes the landmark but does so in an indiscriminate manner, often erasing both the target concept and unrelated contextual elements, which substantially compromises realism.
+ANT is able to achieve more noticeable unlearning, but at the cost of visual artifacts and degraded generation quality, reflecting a lack of fine-grained control over what is erased.
+In contrast, KSCU demonstrates a more controlled behavior: it effectively suppresses the memorized landmark while replacing it with visually plausible alternatives, such as visually similar but semantically distinct objects (e.g., Eiffel Tower $\rightarrow$ highway) or semantically related but non-identical substitutes (e.g., Titanic $\rightarrow$ small boat).
+Such substitutions ensure that the generated images remain coherent and visually compelling, even after the target instance is forgotten.
+Among all evaluated methods, KSCU achieves the most effective and stable unlearning, while simultaneously producing the highest-quality generations.
+These observations highlight that KSCU not only suppresses memorized instances but also substitutes them with plausible alternatives, thereby preserving overall image realism.
+More importantly, KSCU achieves instance unlearning with only 200 parameter updates, and its substantially reduced iterations and lower computational overhead underscore its potential scalability to large-scale unlearning tasks.
+IV-E Ablation Study
+In our ablation study, we primarily evaluated the three major components of KSCU: the Key Step Table, Prompt Augmentation, and Key Step Unlearning Optimization.
+For the Key Step Table ablation, we varied the $S$ and $E$ indices in Algorithm 1, adjusting the step length $L$ to maintain comparable training frequencies.
+This design ensures that performance differences arise primarily from changes in step coverage, rather than training budget.
+By comparing KSCU variants with identical step coverage but different emphasis—such as early 70% vs.
+last 70%, early 80% vs.
+last 80%, and shrink $E$ vs.
+increase $s_{\text{cur}}$ (Table IV, Line 10)—we observe that configurations emphasizing later steps consistently achieve higher unlearning accuracy and robustness.
+This is intuitive: later denoising steps in diffusion sampling capture fine-grained semantic details that directly determine the fidelity of concept representation.
+Hence, unlearning at later steps is inherently more effective.
+As shown in Table IV, configurations covering 100% of steps, the last 80%, and the last 70% achieve comparable unlearning effectiveness.
+However, unlearning accuracy drops by 16.8% when only the last 60% of steps are used.
+These results lead to two key insights: (1) the earliest 30% of steps contribute minimally to the final concept depiction (e.g., nudity) and can be excluded with little impact, thereby saving computation; and (2) steps after the 15th are critical for encoding the target concept and play a dominant role in unlearning effectiveness.
+Another notable observation is that although full-step unlearning achieves slightly higher accuracy, it incurs a larger FID due to broader interference with the model’s generative capacity.
+These findings highlight the efficiency and lower destructiveness of KSCU’s targeted fine-tuning compared to conventional approaches that indiscriminately modify parameters across all denoising steps.
+TABLE IV: Ablation Study: Impact of Key Step Table (with different $S$, $E$, $L$), Key Step Unlearning Optimization, and Prompt Augmentation (PA) on KSCU’s performance.
+[…]
+KSCU early 70% steps 91.1 59.8 15.6 541
+KSCU early 80% steps 92.1 58.1 14.1 624
+KSCU shrink $E$ 88.5 89.3 15.3 570
+KSCU last 60% steps 79.7 85.3 15.7 498
+KSCU last 80% steps 96.1 48.1 14.3 628
+KSCU all steps 96.0 54.5 18.8 782
+KSCU $\lambda_{1}$ = 0 93.6 54.9 13.5 546
+KSCU $\lambda_{2}$ = 0 92.2 57.3 14.1 546
+KSCU w/o PA 92.5 51.8 14.5 549
+KSCU (last 70% steps) 96.5 47.2 14.1 546
+For the Key Step Unlearning Optimization, setting $\lambda_{1}=0$ disables the Unconditional Predicted Noise Optimization component.
+This configuration led to improved retainability but a clear decline in unlearning effectiveness.
+This demonstrates that unconditional noise optimization plays a central role in effective forgetting, albeit at the cost of slightly compromising generation quality.
+Conversely, setting $\lambda_{2}=0$ degrades the optimization objective to preserving the model’s unconditional predicted noise across training.
+This ablation yields noticeably weaker unlearning performance compared to full KSCU, confirming the necessity of balancing both terms.
+Interestingly, the generative retainability of dynamic $\lambda_{2}$ (i.e., standard KSCU) remains comparable to that of $\lambda_{2}=0$, indicating that dynamically weighting $\lambda_{1}$ effectively guides $loss_{\text{regular}}$ to preserve generation in early steps while strengthening unlearning in later ones.
+Finally, for the Prompt Augmentation module, removing it resulted in a clear drop in UA and an increase in ASR, indicating reduced robustness against adversarial or paraphrased prompts.
+Notably, the FID remained almost unchanged, suggesting that Prompt Augmentation imposes minimal impact on generative retainability while substantially enhancing robustness.
+This confirms Prompt Augmentation as a lightweight yet highly effective strategy for improving resilience to real-world prompts without sacrificing image quality.
+IV-F Target Key Step Concept Unlearning
+In Section 3.2, we derived our loss function, which encourages the model to sample in the direction opposite to the original concept $c$.
+This formulation effectively prevents the generation of $c$ by pushing the model toward unrelated alternatives.
+However, the drawback of this approach is that the replacement concept is uncontrolled, often leading to unpredictable or semantically irrelevant outputs.
+To mitigate this limitation, we introduce a slight modification to the loss function that explicitly guides the model to replace $c$ with a user-specified target concept $c^{-}$, rather than leaving the substitution unconstr ained.
+The modified formulation is as follows:
+$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{-}}\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{+}},\end{split}$ (8)
+where $c^{+}$ denotes the concept to be unlearned, and $c^{-}$ is the intended replacement concept.
+Building on this, we define the following modified loss function to explicitly enforce targeted substitution:
+$\text{Loss}_{\text{target unlearn}}=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})\right\|.$ (9)
+This objective provides precise control over the replacement process, ensuring that the model not only forgets $c^{+}$ but also consistently aligns its generations with the specified substitute $c^{-}$.
+Compared to uncontrolled forgetting, this approach transforms concept unlearning into a more structured substitution task, enabling applications such as targeted style transfer or controlled semantic replacement within diffusion models.
+Figure 7 illustrates this behavior, where KSCU successfully replaces bird with balloon and Van Gogh with sketch, confirming the effectiveness of the proposed target unlearning strategy.
+Figure 7: Visualization results of KSCU with target replacement.
+The left column shows generations from SDv1.5, while the right column illustrates the results after applying KSCU.
+Here, the target concepts bird and Van Gogh are replaced with balloon and sketch, respectively.
+V Conclusion
+This study explores the role of denoising steps in concept unlearning and proposes KSCU, which selectively targets key steps for improved effectiveness.
+Experiments across four tasks demonstrate that KSCU outperforms previous SOTA methods in overall performance.
+The key-step perspective introduced here offers valuable insights for future research.
 Appendix A Analysis of Frequency Component Destruction in Diffusion Models
@@
 Since the reverse process restores information corresponding to the degradation in the forward process, proving that diffusion models first sample low-frequency components and later refine high-frequency details requires showing that high-frequency components are the first to be destroyed in the forward process.
-A.1 Definition of the Diffusion Process
+A-A Definition of the Diffusion Process
 We consider a continuous-time diffusion process, where the initial data $\mathbf{x}_{0}$ evolves into $\mathbf{x}_{t}$ over time $t\in[0,T]$.
@@
 Under these assumptions, the simplified diffusion process is given by:
-$d\mathbf{x}_{t}=g(t)d\mathbf{w}_{t},\quad\mathbf{x}_{0}\text{ as the initial condition.}$. (9)
-A.2 Frequency-Domain Analysis
+$d\mathbf{x}_{t}=g(t)d\mathbf{w}_{t},\quad\mathbf{x}_{0}\text{ as the initial condition.}$ (11)
+A-B Frequency-Domain Analysis
 To analyze how different frequency components evolve during diffusion, we apply the Fourier transform to convert the signal from the time domain to the frequency domain.
@@
 Since $d\mathbf{w}_{t}$ represents white noise, its power spectral density is constant across frequencies.
-A.3 Derivation of Frequency Component Degradation
-A.3.1 Representation of the Initial Signal and Noise
+A-C Derivation of Frequency Component Degradation
+A-C1 Representation of the Initial Signal and Noise
 During the forward diffusion process, the signal $\mathbf{x}_{t}$ can be expressed as the sum of the initial signal $\mathbf{x}_{0}$ and accumulated noise:
-[…]
+$\mathbf{x}_{t}=\mathbf{x}_{0}+\int_{0}^{t}g(s)d\mathbf{w}_{s}.$ (12)
 In the frequency domain, the Fourier transform of $\mathbf{x}_{t}$ is given by:
@@
 Since $d\mathbf{w}_{t}$ is white noise, its power spectral density remains uniform across all frequencies.
-A.3.2 Power Spectral Density of the Signal
+A-C2 Power Spectral Density of the Signal
 The power spectral density of $\mathbf{x}_{t}$ is given by:
@@
 Since $\mathbb{E}[\hat{d\mathbf{w}}_{s}(\omega)]=0$, the cross-term vanishes, leaving:
-[…]
-A.3.3 Signal-to-Noise Ratio (SNR) Analysis
+$\mathbb{E}[|\hat{\mathbf{x}}_{t}(\omega)|^{2}]=|\hat{\mathbf{x}}_{0}(\omega)|^{2}+\int_{0}^{t}|g(s)|^{2}ds.$ (16)
+A-C3 Signal-to-Noise Ratio (SNR) Analysis
 The signal-to-noise ratio (SNR) is defined as:
-[…]
+$\text{SNR}(\omega,t)=\frac{|\hat{\mathbf{x}}_{0}(\omega)|^{2}}{\int_{0}^{t}|g(s)|^{2}ds}.$ (17)
 Since noise power is independent of frequency, SNR evolution is dictated entirely by the initial power spectral density $|\hat{\mathbf{x}}_{0}(\omega)|^{2}$.
-A.3.4 Frequency Dependence of SNR Decay
+A-C4 Frequency Dependence of SNR Decay
 For natural signals (e.g., images and audio), the power spectrum typically follows a power-law decay:
@@
 As $\omega$ increases, $|\hat{\mathbf{x}}_{0}(\omega)|^{2}$ rapidly decreases, implying that high-frequency components exhibit lower SNR.
-[…]
-[…]
+Since $\int_{0}^{t}|g(s)|^{2}ds$ increases with time, SNR at all frequencies declines.
+However, because high-frequency components have lower initial power, their SNR reaches a degradation threshold $\text{SNR}_{\text{th}}$ faster than low-frequency components.
+Specifically, the time at which SNR falls below $\text{SNR}_{\text{th}}$ satisfies:
+$\int_{0}^{t_{\text{th}}}|g(s)|^{2}ds=\frac{|\hat{\mathbf{x}}_{0}(\omega)|^{2}}{\text{SNR}_{\text{th}}}.$ (19)
 Since $|\hat{\mathbf{x}}_{0}(\omega)|^{2}$ decreases with $\omega$, $t_{\text{th}}$ is smaller for higher frequencies, meaning high-frequency information is lost earlier.
-A.4 Conclusion
-The analysis confirms that, in the forward diffusion process, high-frequency components are degraded first, followed by lower-frequency components.
-Consequently, in the reverse denoising process, diffusion models reconstruct low-frequency structures first before refining high-frequency details.
-This frequency-dependent degradation provides theoretical justification for the progressive nature of generative sampling in diffusion models.
 Appendix B More Experimental Details
@@
 The visualization results for Stable Diffusion version 1.5 and later versions are presented in the subsequent sections.
-Appendix C Target Concept Unlearning
-In Section 3.2, we derived our loss function, which ensures that the model samples in the direction opposite to the original concept $c$.
-This approach effectively prevents the generation of $c$ by replacing it with an unrelated concept.
-However, the replacement concept is uncontrolled, leading to unpredictable visual outputs.
-To address this, we introduce a slight modification to the loss function, guiding the model to replace $c$ with a specified target concept $c^{-}$, rather than an arbitrary alternative.
-The modified formulation is as follows:
-$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{-}}\\ &=\epsilon_{\theta}(\mathbf{x}_{t},t)+w\cdot\underbrace{(\epsilon^{*}_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t))}_{\text{Guiding towards }c^{+}}\end{split}$ (18)
-where $c^{+}$ represents the concept to be unlearned, and $c^{-}$ denotes the target replacement concept.
-By applying the following modified loss function, we explicitly enforce the substitution of $c^{+}$ with $c^{-}$:
-$\text{Loss}_{\text{target unlearn}}=\left\|\epsilon_{\theta}^{*}(\mathbf{x}_{t},t,\mathbf{c^{+}})-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c^{-}})\right\|.$ (19)
-This formulation allows precise control over the replacement concept, ensuring that the model not only forgets $c^{+}$ but also consistently generates the desired substitute $c^{-}$, thereby enabling targeted concept substitution in diffusion models.
-In the Figure 8, we present the results of KSCU’s target concept unlearning.
-Appendix D Multi-concept Unlearning
-To evaluate the effectiveness of KSCU in multi-concept unlearning, we provide visualized results 7. Our observations indicate that KSCU maintains strong performance in removing multiple concepts.
+Appendix C Multi-concept Unlearning
+To evaluate the effectiveness of KSCU in multi-concept unlearning, we provide visualized results 8. Our observations indicate that KSCU maintains strong performance in removing multiple concepts.
 Figure 8: Results of multi-concept unlearn.
-Appendix E More Visualization Results
-Figure 8: Visualization results of KSCU with target replacement.
-The left side is SDv1.5, and the right side is the result of KSCU.
-The target concepts are bird and Van Gogh, and the replacement concepts are balloon and sketch.
-Figure 9: KSCU’s visualization results for “nudity”, with SDv2.1 on the left and KSCU on the right.
-Figure 10: visualization results of KSCU about artist styles, SDv2.1 on the left and KSCU on the right.
-Figure 11: Visualization results of artist style erasing, model is SDv1.5